### v1

# Generalized finite and affine $W$-algebras in type $A$

Abstract
We construct a new family of affine $W$-algebras $W^{k}(\lambda,\mu)$ parameterized by partitions $\lambda$ and $\mu$ associated with the centralizers of nilpotent elements in $\mathfrak{gl}_{N}$. The new family unifies a few known classes of $W$-algebras. In particular, for the column-partition $\lambda$ we recover the affine $W$-algebras $W^{k}(\mathfrak{gl}_{N},f)$ of Kac, Roan and Wakimoto, associated with nilpotent elements $f\in\mathfrak{gl}_{N}$ of type $\mu$. Our construction is based on a version of the BRST complex of the quantum Drinfeld–Sokolov reduction. We show that the application of the Zhu functor to the vertex algebras $W^{k}(\lambda,\mu)$ yields a family of generalized finite $W$-algebras $U(\lambda,\mu)$ which we also describe independently as associative algebras.
1 Introduction
It is well-known that some non-semisimple Lie algebras can share certain classical properties of their semisimple counterparts. Amongst such examples are centralizers $\mathfrak{a}$ of nilpotent elements in simple Lie algebras. In particular, the remarkable Premet conjecture states that the subalgebra of $\mathfrak{a}$-invariants in the symmetric algebra $S(\mathfrak{a})$ is a free polynomial algebra; see [16]. Although it does not hold in full generality [19], the conjecture inspired further research into these Lie algebras and associated objects. The centers of the universal enveloping algebras $U(\mathfrak{a})$ in type $A$ were constructed explicitly in [4] and generators of the classical $W$-algebras associated with the centralizers were produced in [15].
In the context of the vertex algebra theory, affine $W$-algebras were used in [2] to describe the centers of the universal affine vertex algebras at the critical level associated with the centralizers $\mathfrak{a}$. This description was further applied to quantize the Mishchenko–Fomenko subalgebras in $U(\mathfrak{a})$. The center at the critical level in type $A$ was further investigated in [13], where its explicit generators were produced. It was shown in [14] that analogues of the affine $W$-algebras can be associated with the underlying Lie algebras $\mathfrak{a}$ and their description in terms of generators was also given.
Our goal in this paper is to apply a version of the BRST complex $C^{k}(\lambda,\mu)$ of the quantum Drinfeld–Sokolov reduction to construct a new family of affine $W$-algebras $W^{k}(\lambda,\mu)$ associated with the centralizers $\mathfrak{a}$ of nilpotent elements in the Lie algebra $\mathfrak{gl}_{N}$. Here $\lambda$ is a partition of $N$ with $n$ parts corresponding to the chosen nilpotent element, while $\mu$ is a partition of $n$. We show that $W^{k}(\lambda,\mu)$ inherits a vertex algebra structure from $C^{k}(\lambda,\mu)$ (Theorem 3.5). The main results concerning the structure of $W^{k}(\lambda,\mu)$ are given in Theorem 3.10 and Corollary 3.11, where its generating sets were described.
The family of affine $W$-algebras $W^{k}(\lambda,\mu)$ turns out to interpolate between several classes of vertex algebras previously studied in the literature. In the specialization where $\lambda=(1^{N})$ is the column-partition of $N$ and $\mu=(N)$ is the row-partition, we get the affine $W$-algebra $W^{\hskip 1.0ptk}(\mathfrak{gl}_{N})$ going back to the pioneering works of Zamolodchikov [20] and Fateev and Lukyanov [7]. Such $W$-algebras associated with arbitrary simple Lie algebras originate in conformal field theory and they were intensively studied both in mathematics and physics literature; see e.g. [1], [3] and [8, Ch. 15] for detailed expositions and reviews.
Furthermore, if $\lambda$ is the column-partition while the partition $\mu$ of $N$ is arbitrary, then $W^{k}(\lambda,\mu)$ coincides with the affine $W$-algebra $W^{k}(\mathfrak{gl}_{N},f)$ associated with a nilpotent element $f\in\mathfrak{g}$ of type $\mu$, as introduced by Kac, Roan and Wakimoto [11]; see Remark 3.12 below.
In a different specialization where the partition $\lambda$ with $n$ parts is arbitrary and $\mu=(n)$ is the row-partition, $W^{k}(\lambda,(n))$ is the affine $W$-algebra $W^{k}(\mathfrak{a})$ introduced in a recent work of the second author [14]. For an alternative choice, where $\mu=(1^{n})$ is the column-partition, the $W$-algebra $W^{k}(\lambda,(1^{n}))$ coincides with the universal affine vertex algebra $V^{k}(\mathfrak{a})$, where $\mathfrak{a}$ is the centralizer of a nilpotent element of type $\lambda$. Note that, in general, the $W$-algebra $W^{k}(\lambda,\mu)$ need not be conformal; see Example 5.2.
Using the same partitions $\lambda$ and $\mu$ as above, we also introduce a family of associative algebras $U(\lambda,\mu)$ which we call the generalized finite $W$-algebras; see Definition 2.1. Our main result concerning these objects is Theorem 4.6 which connects the affine $W$-algebra $W^{k}(\lambda,\mu)$ with $U(\lambda,\mu)$ via the Zhu functor originated in [21]. It was already established by De Sole and Kac [6], that the application of the Zhu functor to the affine $W$-algebra $W^{\hskip 1.0ptk}(\mathfrak{gl}_{N},f)$ associated with a nilpotent element $f\in\mathfrak{gl}_{N}$ of type $\mu$ yields the finite $W$-algebra whose definition goes back to Kostant [12]; see [9] and [17]. In our notation, it coincides with $U(\lambda,\mu)$ with $\lambda=(1^{N})$; see Example 2.2 below. Our Theorem 4.6 dealing with arbitrary $\lambda$ and $\mu$ thus provides a generalization of this property and is based on arguments similar to [6]. Note that the finite $W$-algebras $U(\lambda,\mu)$ with $\lambda=(1^{N})$ were also described from a different viewpoint in [5] (denoted there by $W(\pi)$ for the pyramid $\pi$ associated with $\mu$) by using the shifted Yangians; the particular case of rectangular pyramids appeared earlier in [18].
As with the affine case, the algebras $U(\lambda,\mu)$ interpolate between several well-studied objects. The specialization with $\mu=(1^{n})$ and arbitrary $\lambda$ recovers the universal enveloping algebra ${\rm U}(\mathfrak{a})$ of the centralizer $\mathfrak{a}$ of a nilpotent element of type $\lambda$ (Example 2.3). The specialization where $\mu=(n)$ (which we refer to as the ‘principal nilpotent case’) is considered in Theorem 5.4, where we obtain an explicit description of generators of $U(\lambda,(n))$, then show that this algebra is isomorphic to the center of ${\rm U}(\mathfrak{a})$ (Corollary 5.6).
The very last Section 5.2 is devoted to the ‘minimal nilpotent case’ with $\mu=(1^{n-2}2)$ and arbitrary $\lambda$. We give explicit descriptions of generators of both the finite and affine $W$-algebras $U(\lambda,\mu)$ and $W^{k}(\lambda,\mu)$; see the respective Theorems 5.7 and 5.8.
2 Generalized finite $W$-algebras
Throughout the paper $N$ denotes a positive integer and $\mathfrak{g}=\mathfrak{gl}_{N}$ is the general linear Lie algebra over the field $\mathbb{C}$ of complex numbers. Take a nilpotent element $e\in\mathfrak{g}$ whose Jordan canonical form has the Jordan blocks of sizes $\lambda_{1}\leqslant\lambda_{2}\leqslant\dots\leqslant\lambda_{n}.$ Consider the left-justified pyramid corresponding to the partition $\lambda=(\lambda_{1},\dots,\lambda_{n})$ of $N$. It consists of $n$ rows of unit boxes with $\lambda_{1}$ boxes in the top row, $\lambda_{2}$ boxes in the second row, etc. The associated tableau is obtained by writing the numbers $1,2,\dots,N$ into the boxes of the pyramid $\lambda$ consecutively by rows from left to right starting from the top row. For example, the tableau associated with the partition $\lambda=(2,3,5)\vdash 10=N$ is given by
{ytableau}$\ytableau 1&2\\ 345\\ 678910$ (2.1)
We let $\text{row}_{\lambda}(a)$ and $\text{col}_{\lambda}(a)$ denote the row and column number of the box containing the entry $a$. We will denote by $q_{i}$ the number of boxes in the $i$-th column of the pyramid $\lambda$ for $i=1,2,\dots,l,$ where $l:=\lambda_{n}$.
Using the tableau associated with the pyramid $\lambda$, introduce the element $e\in\mathfrak{g}$ by
$e=\sum_{\begin{subarray}{c}i=1,\dots,N-1\\[1.39998pt] \text{row}_{\lambda}(i)=\text{row}_{\lambda}(i+1)\end{subarray}}e_{i\,i+1},$ (2.2)
where the $e_{ij}$ denote the standard basis elements of $\mathfrak{g}$. In the above example, the parameters are $(q_{1},q_{2},q_{3},q_{4},q_{5})=(3,3,2,1,1)$ and the corresponding nilpotent element is
[…]
Given a pyramid $\lambda$ with $n$ rows, take an arbitrary partition $\mu$ of $n$. As with $\lambda$, we will write the parts of $\mu=(\mu_{1},\mu_{2},\dots,\mu_{m})$ in the weakly increasing order $\mu_{1}\leqslant\mu_{2}\leqslant\dots\leqslant\mu_{m}$ and consider the corresponding pyramid $\mu$ along with the associated tableau. For the pyramid $\lambda$ appearing in (2.1), there are three possible pyramids $\mu$ and associated tableaux:
[…]
The centralizer $\mathfrak{a}:=\mathfrak{g}^{e}$ of the element $e\in\mathfrak{g}$ defined in (2.2) is a Lie subalgebra with the basis
$\mathcal{B}^{e}:=\{\ E_{ij}^{(r)}\ |\ (i,j,r)\in S^{e}\,\},$ (2.4)
where
$S^{e}:=\{\,(i,j,r)\in\mathbb{Z}^{3}\,|\,1\leqslant i,j\leqslant n\,,\,\ \lambda_{j}-\text{min}(\lambda_{i},\lambda_{j})\leqslant r<\lambda_{j}\,\}$
and
$E_{ij}^{(r)}=\sum_{\begin{subarray}{c}\text{row}_{\lambda}(a)=i,\,\text{row}_{\lambda}(b)=j\\[1.39998pt] \text{col}_{\lambda}(b)-\text{col}_{\lambda}(a)=r\end{subarray}}e_{ab}.$ (2.5)
The commutation relations in the centralizer $\mathfrak{a}$ are given by
$[E_{ij}^{(r)},E_{hl}^{(s)}]=\delta_{hj}E_{il}^{(r+s)}-\delta_{il}E_{hj}^{(r+s)},$
where we set $E_{ij}^{(r)}=0$ for $(i,j,r)\not\in S^{e}.$
Now we use the chosen pyramid $\mu$ to introduce a $\mathbb{Z}$-gradation $\mathfrak{a}:=\bigoplus_{i\in\mathbb{Z}}\mathfrak{a}(i)$ on the centralizer by setting
$\text{deg}_{\mu}(E_{ij}^{(r)}):=\text{col}_{\mu}(j)-\text{col}_{\mu}(i).$ (2.6)
Set
$\mathfrak{n}_{\lambda,\mu}\quad:=\quad\bigoplus_{i>0}\mathfrak{a}(i)\quad=\quad\text{Span}_{\mathbb{C}}\ \{\ E_{ij}^{(r)}\ |\ (i,j,r)\in S_{\lambda,\mu}\},$ (2.7)
where
$S_{\lambda,\mu}:=\{(i,j,r)\in S^{e}\,|\,\text{col}_{\mu}(i)<\text{col}_{\mu}(j)\}.$ (2.8)
Introduce the element
$\mathsf{e}:=\sum_{\begin{subarray}{c}i=1,\dots,n-1\\[1.39998pt] \text{row}_{\mu}(i)=\text{row}_{\mu}(i+1)\end{subarray}}E_{i\,i+1}^{(\lambda_{i+1}-1)}\in\mathfrak{a}(1)$ (2.9)
associated with the pair of pyramids $(\lambda,\mu)$. To illustrate, note that the elements associated with the respective tableaux in (2.3) are
$\mathsf{e}=0,\qquad\mathsf{e}=E_{23}^{(4)}\qquad\text{and}\qquad\mathsf{e}=E_{12}^{(2)}+E_{23}^{(4)}.$
Obviously, in the general case the relation $\text{row}_{\mu}(i)=\text{row}_{\mu}(i+1)$ is possible if and only if $\text{col}_{\mu}(i+1)=\text{col}_{\mu}(i)+1$. Define the associated character $\chi\in\mathfrak{n}_{\lambda,\mu}^{\ast}$ by setting
$\chi\big{(}E_{i\,i+1}^{(\lambda_{i+1}-1)})=1$ (2.10)
if $\text{row}_{\mu}(i)=\text{row}_{\mu}(i+1)$ with $i\in\{1,\dots,n-1\}$, and $\chi(E_{ij}^{(r)})=0$ for all remaining triples $(i,j,r)\in S_{\lambda,\mu}$. Consider the left ideal
$\mathcal{I}_{\lambda,\mu}:=U(\mathfrak{a})\left<\left.\mathsf{n}+\chi(\mathsf{n})\right|\mathsf{n}\in\mathfrak{n}_{\lambda,\mu}\,\right>$
of the universal enveloping algebra $U(\mathfrak{a})$. It is clear that the adjoint action of $\mathfrak{n}_{\lambda,\mu}$ on $U(\mathfrak{a})$ induces the action of $\mathfrak{n}_{\lambda,\mu}$ on the quotient $U(\mathfrak{a})/\mathcal{I}_{\lambda,\mu}$. As with the finite $W$-algebras (cf. [5] and [6, Appendix]), the subspace of $\mathfrak{n}_{\lambda,\mu}$-invariants turns out to be an associative algebra with the product inherited from $U(\mathfrak{a})$ as we verify below.
Definition 2.1.
The generalized finite $W$-algebra is the associative algebra
$U(\lambda,\mu)=(U(\mathfrak{a})/\mathcal{I}_{\lambda,\mu})^{\mathfrak{n}_{\lambda,\mu}}.$
To verify that $U(\lambda,\mu)$ is a well-defined associative algebra, take $A,B\in U(\mathfrak{a})$ such that $\overline{A},\overline{B}\in U(\lambda,\mu)$ are their images in the quotient, and $\mathsf{n}\in\mathfrak{n}_{\lambda,\mu}$. It is sufficient to show that $[\mathsf{n},AB]\in\mathcal{I}_{\lambda,\mu}$. Since $\overline{A}\in U(\lambda,\mu)$, we have $[\mathsf{n},A]=a(\mathsf{m}+\chi(\mathsf{m}))$ for some $a\in U(\mathfrak{a})$ and $\mathsf{m}\in\mathfrak{n}_{\lambda,\mu}$. By the Leibniz rule,
$[\mathsf{n},AB]$ $=[\mathsf{n},A]B+A[\mathsf{n},B]$
$=a(\mathsf{m}+\chi(\mathsf{m}))B+A[\mathsf{n},B]$
$=aB(\mathsf{m}+\chi(\mathsf{m}))+a[\mathsf{m},B]+A[\mathsf{n},B]\in\mathcal{I}_{\lambda,\mu},$
as required.
Example 2.2.
Let $\lambda=(1,1,\dots,1)$ be the column-pyramid with $N$ boxes so that $\mathfrak{a}=\mathfrak{gl}_{N}$ and let $\mu$ be a pyramid with $N$ boxes. Consider the element $\mathsf{e}$ associated with $\mu$ as defined in (2.9). Then
$U(\lambda,\mu)=U(\mathfrak{g},\mathsf{e})$
is the finite $W$-algebra associated with $\mathfrak{g}$ and $\mathsf{e}$; see [9], [12] and [17].
Example 2.3.
Let $\lambda$ be a pyramid consisting of $N$ boxes with $n$ rows and $\mu=(1,1,\dots,1)$ be the column-pyramid with $n$ boxes. Then
$U(\lambda,\mu)=U(\mathfrak{a}),$
where $\mathfrak{a}$ is the centralizer of the nilpotent element $e$ defined in (2.2).
3 Generalized affine $W$-algebras
3.1 Affine vertex algebra and free fermion vertex algebra
Let $e$ be the nilpotent element in $\mathfrak{g}$ associated with the pyramid $\lambda$ as defined in (2.2). As before, we denote by $\mathfrak{a}$ be the centralizer of $e$ in $\mathfrak{g}.$ We will use the invariant bilinear form (|)(\ |\ )( | ) on $\mathfrak{a}$ which was introduced in [2]. Specifically, set $\text{deg}_{\lambda}(e_{ij})=\text{col}_{\lambda}(j)-\text{col}_{\lambda}(i)$ to induce the $\mathbb{Z}$-gradations
$\mathfrak{g}=\bigoplus_{r\in\mathbb{Z}}\mathfrak{g}_{r}\qquad\text{and}\qquad\mathfrak{a}=\bigoplus_{r\in\mathbb{Z}}\mathfrak{a}_{r}=\bigoplus_{r\in\mathbb{Z}}(\mathfrak{a}\cap\mathfrak{g}_{r}).$
Then for homogeneous elements $X,Y\in\mathfrak{a}$ we set
$(X|Y)=\begin{cases}\frac{1}{2N}\text{tr}_{\mathfrak{g}_{0}}\big{(}(\text{ad }X)(\text{ad }Y)\big{)}\quad&\text{if }X,Y\in\mathfrak{a}_{0},\\ 0\quad&\text{otherwise.}\end{cases}$ (3.1)
To use a version of the BRST complex of the quantum Drinfeld–Sokolov reduction, associated with the Lie algebra $\mathfrak{a}$ by analogy with [8, Ch. 15], and extending the construction of [14], introduce two vertex algebras as follows. The first is $V^{k}(\mathfrak{a})$, the affine vertex algebra at the level $k\in\mathbb{C}$. It is defined as
[…]
$[a\otimes t^{n},b\otimes t^{m}]=[a,b]\otimes t^{n+m}+\delta_{n+m,0}(a|b)K$ (3.2)
using the bilinear form (3.1). Here $\mathbb{C}_{k}$ is the one-dimensional module of $U(\mathfrak{a}[t])\oplus\mathbb{C}K$, where all elements act trivially except $K$, which acts as $k\in\mathbb{C}$. For $(E_{ij}^{(r)}\otimes t^{-1})\in V^{k}(\mathfrak{a})$, the corresponding field is $E_{ij}^{(r)}(z):=\sum_{n\in\mathbb{Z}}(E_{ij}^{(r)}\otimes t^{n})z^{-n-1}$ and the commutation relations induced from (3.2) are
$[E_{i_{1},j_{1}}^{(r_{1})}(z),E_{i_{2},j_{2}}^{(r_{2})}(w)]=[E_{i_{1},j_{1}}^{(r_{1})},E_{i_{2},j_{2}}^{(r_{2})}](w)\delta(z,w)+k(E_{i_{1},j_{1}}^{(r_{1})}\ |\ E_{i_{2},j_{2}}^{(r_{2})})\partial_{w}\delta(z,w),$
where $\delta(z,w)=\sum_{m\in\mathbb{Z}}z^{m}w^{-m-1}.$ In terms of the $\lambda$-bracket, they can be rewritten as
$[E_{i_{1},j_{1}}^{(r_{1})}{}_{\lambda}E_{i_{2},j_{2}}^{(r_{2})}]=[E_{i_{1},j_{1}}^{(r_{1})},E_{i_{2},j_{2}}^{(r_{2})}]+k(E_{i_{1},j_{1}}^{(r_{1})}\ |\ E_{i_{2},j_{2}}^{(r_{2})})\,\lambda;$
see e.g. [10].
The second vertex algebra is $\mathcal{F}(\mathfrak{n}_{\lambda,\mu})$, the free fermion vertex algebra associated with $\mathfrak{n}_{\lambda,\mu}$. More precisely, let us consider the dual space $\mathfrak{n}_{\lambda,\mu}^{\ast}$ of $\mathfrak{n}_{\lambda,\mu}$ and the odd vector superspaces
$\phi_{\mathfrak{n}_{\lambda,\mu}}=\{\phi_{\mathsf{n}}\,|\,\mathsf{n}\in\mathfrak{n}_{\lambda,\mu}\}\qquad\text{and}\qquad\phi^{\mathfrak{n}_{\lambda,\mu}^{\ast}}=\{\phi^{\mathsf{m}}\,|\,\mathsf{m}\in\mathfrak{n}_{\lambda,\mu}^{\ast}\}.$ (3.3)
We set for brevity,
$\phi_{(i,j,r)}:=\phi_{E_{ij}^{(r)}}\qquad\text{and}\qquad\phi^{(i,j,r)}:=\phi^{E_{ij}^{(r)\ast}},$
where the $E_{ij}^{(r)\ast}$ are the elements of the basis of $\mathfrak{n}_{\lambda,\mu}^{\ast}$ dual to the basis formed by the elements $E_{ij}^{(r)}\in\mathfrak{n}_{\lambda,\mu}$. The vertex algebra $\mathcal{F}(\mathfrak{n}_{\lambda,\mu})$ is freely generated by the elements $\phi_{(i,j,r)}$ and $\phi^{(i,j,r)}$ for $(i,j,r)\in S_{\lambda,\mu}$ as a differential algebra with the following $\lambda$-brackets:
$[\phi_{(i,j,r)}{}_{\lambda}\phi^{(i^{\prime},j^{\prime},r^{\prime})}]$ $=[\phi^{(i^{\prime},j^{\prime},r^{\prime})}{}_{\lambda}\phi_{(i,j,r)}]=\delta_{ii^{\prime}}\delta_{jj^{\prime}}\delta_{rr^{\prime}},$ (3.4)
$[\phi^{(i,j,r)}{}_{\lambda}\phi^{(i^{\prime},j^{\prime},r^{\prime})}]$ $=[\phi_{(i,j,r)}{}_{\lambda}\phi_{(i^{\prime},j^{\prime},r^{\prime})}]=0.$
The first relation in (3.4) can also be written as
$[\phi_{\mathsf{n}}{}_{\lambda}\phi^{\mathsf{m}}]=[\phi^{\mathsf{m}}{}_{\lambda}\phi_{\mathsf{n}}]={\mathsf{m}}({\mathsf{n}})$
for ${\mathsf{n}}\in\mathfrak{n}_{\lambda,\mu}$ and ${\mathsf{m}}\in\mathfrak{n}_{\lambda,\mu}^{\ast}.$ Also, we set $\phi_{x}:=\phi_{\pi_{+}(x)}$ where $\pi_{+}:\mathfrak{a}\to\mathfrak{n}_{\lambda,\mu}$ is the projection to $\mathfrak{n}_{\lambda,\mu}$ which is zero on the basis vectors which do not belong to $\mathfrak{n}_{\lambda,\mu}$.
3.2 Definition of the generalized affine $W$-algebras
In this section, we introduce the generalized affine $W$-algebras via BRST cohomologies. Consider the vertex algebra
$C^{k}(\lambda,\mu):=V^{k}(\mathfrak{a})\otimes\mathcal{F}(\mathfrak{n}_{\lambda,\mu})$
and its element
$d=\sum_{I\in S_{\lambda,\mu}}:\phi^{I}E_{I}:+\phi^{\chi}+\frac{1}{2}\sum_{I,I^{\prime}\in S_{\lambda,\mu}}:\phi^{I}\phi^{I^{\prime}}\phi_{[E_{I^{\prime}},E_{I}]}:\ \in\ C^{k}(\lambda,\mu).$
Here $E_{I}:=E_{ij}^{(r)}$ and $\phi^{I}:=\phi^{(i,j,r)}$ for $I=(i,j,r)\in S_{\lambda,\mu}$; see also (2.10) for the definition of $\chi$. Let us denote
$Q:=d_{(0)}\ \in\ \text{End}\,C^{k}(\lambda,\mu),$
where $d_{(0)}:A\mapsto[d\,{}_{\lambda}\,A]\,\big{|}_{\lambda=0}.$ By the fundamental property of $\lambda$-bracket, $Q$ is an odd derivation with respect to the normally ordered product and commutes with $\partial$. Hence the following lemma completely determines $Q.$
Lemma 3.1.
The following formulas hold:
  1. (1)
    $Q(A)=\sum_{I\in S_{\lambda,\mu}}:\phi^{I}[E_{I},a]:+\sum_{(i,j,r)\in S_{\lambda,\mu}}k(a|E_{I})\partial\phi^{I}$ for $a\in\mathfrak{a}$.
  2. (2)
    $Q(\phi_{\mathsf{n}})=\mathsf{n}+\chi(\mathsf{n})+\sum_{I\in S_{\lambda,\mu}}:\phi^{I}\phi_{[E_{I},\mathsf{n}]}:$ for ${\mathsf{n}}\in\mathfrak{n}_{\lambda,\mu}$.
  3. (3)
    $Q(\phi^{\mathsf{m}})=\frac{1}{2}\sum_{I\in S_{\lambda,\mu}}:\phi^{I}\phi^{E_{I}\cdot\mathsf{m}}:$ for ${\mathsf{m}}\in\mathfrak{n}_{\lambda,\mu}^{\ast}$.
In the last equation, $E_{I}\cdot\mathsf{m}$ denotes the coadjoint action of $\mathfrak{n}_{\lambda,\mu}$ on $\mathfrak{n}_{\lambda,\mu}^{\ast}$.
Proof.
It is enough to compute $\lambda$-brackets between $d$ and generators of $C^{k}(\lambda,\mu).$
  1. (1)
    Observe that $[d{}_{\lambda}a]=\sum_{I\in S_{\lambda,\mu}}[\,:\phi^{I}E_{I}:\,{}_{\lambda}\,a\,].$ By skew-symmetry, we have
    $[d{}_{\lambda}a]$ $=-\sum_{I\in S_{\lambda,\mu}}\big{[}a_{-\lambda-\partial}:\phi^{I}E_{I}:\big{]}=\sum_{I\in S_{\lambda,\mu}}:\phi^{I}[E_{I},a]:+\sum_{I\in S_{\lambda,\mu}}k(\partial+\lambda)(a|E_{I})\phi^{I}.$
  2. (2)
    By direct computations, we get $\big{[}\phi^{\chi}\,{}_{\lambda}\ \phi_{\mathsf{n}}\big{]}=\chi(\mathsf{n})$ and
    $\sum_{I\in S_{\lambda,\mu}}\big{[}:\phi^{I}E_{I}:{}_{\lambda}\ \phi_{\mathsf{n}}\big{]}$ $=\sum_{I\in S_{\lambda,\mu}}:[\phi_{\mathsf{n}\ -\lambda-\partial}\ \phi^{I}]\,E_{I}:=\mathsf{n}.$
    We also have
    $\sum_{{I,I^{\prime}}\,\in S_{\lambda,\mu}}\big{[}:\phi^{I}\phi^{I^{\prime}}\phi_{[E_{I^{\prime}},E_{I}]}:\,{}_{\lambda}\phi_{\mathsf{n}}\big{]}$
    $:\displaystyle=\sum_{I,I^{\prime}\in S_{\lambda,\mu}}:[\phi_{\mathsf{n}\ -\lambda-\partial}\phi^{I}]:\phi^{I^{\prime}}\phi_{[E_{I^{\prime}},E_{I}]}::-\sum_{I,I^{\prime}\,\in S_{\lambda,\mu}}:\phi^{I}[\phi_{\mathsf{n}\ -\lambda-\partial}:\phi^{I^{\prime}}\phi_{[E_{I^{\prime}},E_{I}]}:]$ :
    $=\sum_{I^{\prime}\in S_{\lambda,\mu}}:\phi^{I^{\prime}}\phi_{[E_{I^{\prime}},\mathsf{n}]}:-\sum_{I\in S_{\lambda,\mu}}:\phi^{I}\phi_{[\mathsf{n},E_{I}]}:\ =2\sum_{(i,j,r)\in S_{\lambda,\mu}}:\phi^{I}\phi_{[E_{I},\mathsf{n}]}:\ .$
    Hence we conclude that
    $[d_{\lambda}\phi_{\mathsf{n}}]=\mathsf{n}+\chi(\mathsf{n})+\sum_{I\in S_{\lambda,\mu}}:\phi^{I}\phi_{[E_{I},\mathsf{n}]}:\ .$
  3. (3)
    We have
    $[d{}_{\lambda}\phi^{\mathsf{m}}]$ $::\displaystyle=\frac{1}{2}\sum_{I,I^{\prime}\in S_{\lambda,\mu}}:\phi^{I}:\phi^{I^{\prime}}\big{[}\phi^{\mathsf{m}}_{\ \lambda}\phi_{[E_{I^{\prime}},E_{I}]}\big{]}:$
    $=\frac{1}{2}\sum_{I,I^{\prime}\in S_{\lambda,\mu}}\mathsf{m}\left([E_{I^{\prime}},E_{I}]\right):\phi^{I}\phi^{I^{\prime}}:\ =\frac{1}{2}\sum_{I\in S_{\lambda,\mu}}:\phi^{I}\phi^{E_{I}\cdot\mathsf{m}}:\ .$
Proposition 3.2.
The endomorphism $Q$ is a differential on $C^{k}(\lambda,\mu)$; that is, $Q^{2}=0.$
Proof.
To verify the relation $Q^{2}=0$, it is enough to show that $[d{}_{\lambda}d]=0$. We have
$\sum_{I\in S_{\lambda,\mu}}\big{[}\,d\,{}_{\lambda}:\phi^{I}E_{I}:\big{]}=\sum_{I\in S_{\lambda,\mu}}\Big{(}:[d_{\lambda}\phi^{I}]E_{I}:-:\phi^{I}[d_{\lambda}E_{I}]:+\int_{0}^{\lambda}[[d_{\lambda}\phi^{I}]_{\mu}E_{I}]\ d\mu\Big{)}$
$::\displaystyle=\sum_{I,I^{\prime}\in S_{\lambda,\mu}}\frac{1}{2}::\phi^{I^{\prime}}\phi^{E_{I^{\prime}}\cdot E_{I}^{\ast}}:E_{I}:-\sum_{I,I^{\prime}\in S_{\lambda,\mu}}:\phi^{I}:\phi^{I^{\prime}}[E_{I^{\prime}},E_{I}]:$
$::\displaystyle=\sum_{I,I^{\prime}\in S_{\lambda,\mu}}\frac{1}{2}::\phi^{I}\phi^{I^{\prime}}:[E_{I^{\prime}},E_{I}]:-\sum_{I,I^{\prime}\in S_{\lambda,\mu}}:\phi^{I}:\phi^{I^{\prime}}[E_{I^{\prime}},E_{I}]:$
$=-\frac{1}{2}\sum_{I,I^{\prime}\in S_{\lambda,\mu}}:\phi^{I}:\phi^{I^{\prime}}[E_{I^{\prime}},E_{I}]::\ .$
One can easily check that $[d_{\lambda}\phi^{\chi}]=0$ and hence it remains to verify that
$\frac{1}{2}\sum_{I,I^{\prime}\in S_{\lambda,\mu}}\big{[}d_{\lambda}:\phi^{I}\phi^{I^{\prime}}\phi_{[E_{I^{\prime}},E_{I}]}:\big{]}=-\sum_{I\in S_{\lambda,\mu}}\big{[}d_{\lambda}:E_{I}\phi^{I}:\big{]}.$ (3.5)
Expand the left hand side of (3.5) as follows:
$\frac{1}{2}\sum_{I,I^{\prime}\in S_{\lambda,\mu}}\big{[}d_{\lambda}:\phi^{I}\phi^{I^{\prime}}\phi_{[E_{I^{\prime}},E_{I}]}:\big{]}$ (3.6)
$::\displaystyle=\frac{1}{4}\sum_{I,I^{\prime},I^{\prime\prime},I^{\prime\prime\prime}\in S_{\lambda,\mu}}(E_{I})^{\ast}\left([E_{I^{\prime\prime\prime}},E_{I^{\prime\prime}}]\right)::\phi^{I^{\prime\prime}}\phi^{I^{\prime\prime\prime}}:\ :\phi^{I^{\prime}}\phi_{[E_{I^{\prime}},E_{I}]}:$ (3.7)
$-\frac{1}{4}\sum_{I,I^{\prime},I^{\prime\prime},I^{\prime\prime\prime}\in S_{\lambda,\mu}}(E_{I^{\prime}})^{\ast}\left([E_{I^{\prime\prime\prime}},E_{I^{\prime\prime}}]\right):\phi^{I}(::\phi^{I^{\prime\prime}}\phi^{I^{\prime\prime\prime}}:\phi_{[E_{I^{\prime}},E_{I}]}:):$ (3.8)
$+\frac{1}{2}\sum_{I,I^{\prime}\in S_{\lambda,\mu}}:\phi^{I}:\phi^{I^{\prime}}[E_{I^{\prime}},E_{I}]::$ (3.9)
$+\frac{1}{2}\sum_{I,I^{\prime}\in S_{\lambda,\mu}}\chi\left([E_{I^{\prime}},E_{I}]\right):\phi^{I}\phi^{I^{\prime}}:$ (3.10)
$+\frac{1}{2}\sum_{I,I^{\prime},I^{\prime\prime}\in S_{\lambda,\mu}}:\phi^{I}:\phi^{I^{\prime}}:\phi^{I^{\prime\prime}}\phi_{[E_{I^{\prime\prime}},[E_{I^{\prime}},E_{I}]]}:::$ (3.11)
$-\frac{1}{2}\sum_{I,I^{\prime},I^{\prime\prime}\in S_{\lambda,\mu}}E_{I^{\prime}}^{\ast}([[E_{I^{\prime}},E_{I}],E_{I^{\prime\prime}}]):\phi^{I}\phi^{I^{\prime\prime}}:\ \lambda$ (3.12)
$+\frac{1}{2}\sum_{I,I^{\prime},I^{\prime\prime}\in S_{\lambda,\mu}}E_{I^{\prime}}^{\ast}([[E_{I^{\prime\prime}},E_{I^{\prime}}],E_{I}]):\phi^{I}\phi^{I^{\prime\prime}}:\ \lambda\ .$ (3.13)
The expression in (3.10) is zero because $\text{deg}([E_{I^{\prime}},E_{I}])\geqslant 2$ (if the commutator is nonzero). The sum of the terms in (3.12) and (3.13) is zero, as follows by relabeling the indices $I$ and $I^{\prime\prime}$ in (3.13), while the sum of the terms in (3.7), (3.8) and (3.11) is zero because of the Jacobi identity. Hence (3.5) follows. ∎
By Proposition 3.2, the cohomology $H(C^{k}(\lambda,\mu),Q)$ is well-defined. This enables us to state the key definition.
Definition 3.3.
The vertex algebra
$W^{k}(\lambda,\mu)=H(C^{k}(\lambda,\mu),Q)$
is called the generalized affine $W$-algebra of level $k$ associated with the partitions $\lambda$ and $\mu$.
Recall that the operator $Q=d_{(0)}$ commutes with the endomorphism $\partial$. Note also that $Q$ is a derivation with respect to the normally ordered product and $\lambda$-bracket. In other words, we have the properties
$Q(:AB:)=:Q(A)B:+(-1)^{p(A)}:A\,Q(B):$ (3.14)
$Q([A\,{}_{\lambda}\,B])=[Q(A)_{\lambda}B]+(-1)^{p(A)}[A_{\lambda}Q(B)]$ (3.15)
for $A,B\in C^{k}(\lambda,\mu).$ Indeed, (3.14) follows from the Wick formula, whereas (3.15) follows from the Jacobi identity of Lie conformal algebras. These two properties imply that if $A$ and $B$ in $C^{k}(\lambda,\mu)$ are representatives of elements in $W^{k}(\lambda,\mu)$, i.e. $Q(A)=Q(B)=0$, then $Q(:AB:)=Q([A{}_{\lambda}B])=0$. Furthermore, we have the following lemma which implies that a vertex algebra structure on $W^{k}(\lambda,\mu)$ is well-defined.
Lemma 3.4.
  1. (1)
    If $A\in C^{k}(\lambda,\mu)$ is in the image of $Q$ then $\partial A$ is also in the image of $Q$.
  2. (2)
[…]
  3. (3)
    If $B\in\ker Q$ and $A\in C^{k}(\lambda,\mu)$ then $[Q(A)_{\lambda}B]$ is in $\mathbb{C}[\lambda]\otimes\text{im}(Q)$.
Proof.
Suppose $Q(B)=A$ for some $B\in C^{k}(\lambda,\mu).$ Since $Q(\partial B)=\partial A,$ the first assertion follows. The second and third assertions follow from (3.14) and (3.15), respectively, by taking into account the relation $[A_{\lambda}Q(B)]=0.$
Theorem 3.5.
The vertex algebra structure of $C^{k}(\lambda,\mu)$ induces the vertex algebra structure on $W^{k}(\lambda,\mu)$.
Proof.
By Lemma 3.4, $W^{k}(\lambda,\mu)$ is closed under the derivation $\partial$, normally ordered product, and the $\lambda$-bracket. ∎
3.3 Structure of the generalized affine $W$-algebra
In this section, we describe some properties of a minimal generating set of the vertex algebra $W^{k}(\lambda,\mu)$ by introducing another construction via a subcomplex of $C^{k}(\lambda,\mu).$ Introduce the ‘building blocks’
$J_{a}=a+\sum_{I\in S_{\lambda,\mu}}:\phi^{I}\phi_{[E_{I},a]}:$ (3.16)
for $a\in\mathfrak{a}$ and let $\pi_{\leqslant}a:=a-\pi_{+}(a)$. By direct computations, we can check that
$[d_{\lambda}J_{a}]$ $=\sum_{I\in S_{\lambda,\mu}}:\phi^{I}J_{\pi_{\leq}[E_{I},a]}:-\ \phi^{a\cdot\chi}$ (3.17)
$+(\lambda+\partial)\sum_{I\in S_{\lambda,\mu}}\Big{(}k(E_{I}|a)+\text{tr}_{\mathfrak{n}_{\lambda,\mu}}(\pi_{+}\text{ad }E_{I}\circ\pi_{+}\text{ad }a)\Big{)}\phi^{I}$
where $a\cdot\chi$ is the coadjoint action of $\mathfrak{n}_{\lambda,\mu}$ on $\mathfrak{n}_{\lambda,\mu}^{\ast}$. Also, we have
$[J_{a}{}_{\lambda}\phi_{\mathsf{n}}]=\phi_{[a,\mathsf{n}]},\quad[J_{a}{}_{\lambda}\phi^{\mathsf{m}}]=\phi^{a\cdot\mathsf{m}}\ ,$ (3.18)
and
$[J_{a\ \lambda}J_{b}]$ $=J_{[a,b]}+(k(a|b)+\text{tr}_{\mathfrak{n}_{\lambda,\mu}}(\pi_{+}\text{ad }a\circ\pi_{+}\text{ad }b))\lambda$
$-\sum_{I\in S_{\lambda,\mu}}:\phi^{I}\phi_{[\pi_{\leq}[E_{I},a],b]}:+\sum_{I\in S_{\lambda,\mu}}:\phi^{I}\phi_{[\pi_{\leq}[E_{I},b],a]}:\ .$
In particular, if $a$ and $b$ are both in $\mathfrak{n}_{\lambda,\mu}$ or in the subspace
$\mathfrak{p}:=\bigoplus_{i\leqslant 0}\mathfrak{a}(i),$ (3.19)
then we have
$[J_{a}{}_{\lambda}J_{b}]=J_{[a,b]}+(k(a|b)+\text{tr}_{\mathfrak{n}_{\lambda,\mu}}(\pi_{+}\text{ad }a\circ\pi_{+}\text{ad }b))\lambda.$ (3.20)
We have the tensor product decomposition of the vertex algebra $C^{k}(\lambda,\mu)$ given by
$C^{k}(\lambda,\mu)=C_{+}^{k}(\lambda,\mu)\otimes\widetilde{C}^{k}(\lambda,\mu),$
where $C_{+}^{k}(\lambda,\mu)$ and $\widetilde{C}^{k}(\lambda,\mu)$ are the vertex subalgebras respectively generated by the subspaces $r_{+}:=\phi_{\mathfrak{n}_{\lambda,\mu}}\oplus J_{\mathfrak{n}_{\lambda,\mu}}$ and $r_{-}:=\phi^{\mathfrak{n}_{\lambda,\mu}^{\ast}}\oplus J_{\mathfrak{p}}$. By (3.18) and (3.20), we can conclude $C_{+}^{k}(\lambda,\mu)$ and $\widetilde{C}^{k}(\lambda,\mu)$ are universal enveloping vertex algebras of the nonlinear Lie conformal algebras
$R_{+}:=\mathbb{C}[\partial]\otimes r_{+}\qquad\text{and}\qquad R_{-}:=\mathbb{C}[\partial]\otimes r_{-}.$
Moreover, the differential $Q$ has the properties
$Q(C_{+}^{k}(\lambda,\mu))\subset C_{+}^{k}(\lambda,\mu)\qquad\text{and}\qquad Q(\widetilde{C}^{k}(\lambda,\mu))\subset\widetilde{C}^{k}(\lambda,\mu).$
Due to the fact that $H(C_{+}^{k}(\lambda,\mu),Q|_{C_{+}^{k}(\lambda,\mu)})=\mathbb{C}$, we then have
$H(C^{k}(\lambda,\mu),Q)\simeq H(C_{+}^{k}(\lambda,\mu),Q|_{C_{+}^{k}(\lambda,\mu)})\otimes H(\widetilde{C}^{k}(\lambda,\mu),\widetilde{Q})\simeq H(\widetilde{C}^{k}(\lambda,\mu),\widetilde{Q}),$
where $\widetilde{Q}:=Q|_{\widetilde{C}^{k}(\lambda,\mu)}$. As a conclusion, the following proposition holds.
Proposition 3.6.
The generalized $W$-algebra $W^{k}(\lambda,\mu)$ is isomorphic to $H(\widetilde{C}^{k}(\lambda,\mu),\widetilde{Q}).$
Define the conformal weight $\Delta$ on the complex $\widetilde{C}^{k}(\lambda,\mu)$ induced by
$\Delta(J_{a})=1-j_{a},\quad\Delta(\phi^{I})=\text{deg}(E_{I}),\quad\Delta(\partial)=1$ (3.21)
for $a\in\mathfrak{a}(j_{a})$. We denote the subspace of conformal weight $\Delta$ of $\widetilde{C}^{k}(\lambda,\mu)$ by $\widetilde{C}^{k}(\lambda,\mu)[\Delta].$ Observe that the differential $d$ preserves the conformal grading and
$\widetilde{C}^{k}(\lambda,\mu)[\Delta_{1}]\,{}_{(n)}\,\widetilde{C}^{k}(\lambda,\mu)[\Delta_{2}]\ \subset\ \widetilde{C}^{k}(\lambda,\mu)[\Delta_{1}+\Delta_{2}-n-1].$
Note that an endomorphism $H$ of $\widetilde{C}^{k}(\lambda,\mu)$ such that $H(A)=\Delta(A)A$ for any homogeneous element $A\in\widetilde{C}^{k}(\lambda,\mu)$ is a Hamiltonian operator. This observation will be used in Section 4.1.
Remark 3.7.
If a Lie algebra admits a nondegenerate bilinear form, the corresponding affine vertex algebra at a non-critical level contains a conformal vector. This is known as the Sugawara construction, and it induces conformal vectors of the $W$-algebras (off the critical level) which are quantum Hamiltonian reductions of the affine vertex algebras. However, since our bilinear form (3.1) is degenerate, the affine vertex algebra $V^{k}(\mathfrak{a})$ need not have the Sugawara operator, and the conformal weight (3.21) does not come from a conformal vector of $W^{k}(\lambda,\mu)$; see also Example 5.2. ∎
We will need the $\mathbb{Z}/2$-bigrading on $\widetilde{C}^{k}(\lambda,\mu)$ defined as follows:
$\text{gr}(J_{a})$ $=\left(j_{a}-\frac{1}{2},-j_{a}+\frac{1}{2}\right),\qquad\text{gr}(\partial)=(0,0),$ (3.22)
$\text{gr}(\phi^{I})$ $=\left(\text{deg}(E_{I})+\frac{1}{2},-\text{deg}(E_{I})+\frac{1}{2}\right).$
The bigrading induces the $\mathbb{Z}_{+}$-grading on $\widetilde{C}^{k}(\lambda,\mu)$ given by
$\widetilde{C}^{k}(\lambda,\mu):=\bigoplus_{n\in\mathbb{Z}_{+}}\widetilde{C}^{k}(\lambda,\mu)^{n},$
where
$\widetilde{C}^{k}(\lambda,\mu)^{n}=\text{Span}_{\mathbb{C}}\big{\{}\,:a_{1}a_{2}\dots a_{s}:\,|\,a_{k}\in R_{-},\ \text{gr}(a_{k})=(p_{k},q_{k}),\ \sum_{k=1}^{s}(p_{k}+q_{k})=n\,\big{\}}.$
It also induces the decreasing filtration
$F^{p}(\widetilde{C}^{k}(\lambda,\mu))=\text{Span}_{\mathbb{C}}\big{\{}\,:a_{1}a_{2}\dots a_{s}:\,|\,\,a_{k}\in R_{-},\ \text{gr}(a_{k})=(p_{k},q_{k}),\ \sum_{k=1}^{s}p_{k}\geqslant p\,\big{\}}.$ (3.23)
One easily checks that
$\widetilde{Q}(F^{p}\widetilde{C}^{k}(\lambda,\mu)^{n}[\Delta])\subset F^{p}\widetilde{C}^{k}(\lambda,\mu)^{n+1}[\Delta].$
The associated graded algebra is then defined by
$\text{gr}(\widetilde{C}^{k}(\lambda,\mu)):=\bigoplus_{p,q\in\mathbb{Z}/{2}}\text{gr}^{pq}\widetilde{C}^{k}(\lambda,\mu),$ (3.24)
where
$\text{gr}^{pq}\widetilde{C}^{k}(\lambda,\mu)=F^{p}(\widetilde{C}^{k}(\lambda,\mu)^{p+q})/F^{p+\frac{1}{2}}(\widetilde{C}^{k}(\lambda,\mu)^{p+q})$
and
$F^{p}(\widetilde{C}^{k}(\lambda,\mu)^{p+q})=F^{p}(\widetilde{C}^{k}(\lambda,\mu))\cap\widetilde{C}^{k}(\lambda,\mu)^{p+q}.$
The corresponding graded cohomology is defined by
$H^{pq}(\text{gr}(\widetilde{C}^{k}(\lambda,\mu)),\widetilde{Q}^{\text{gr}})=\frac{\text{ker}(\widetilde{Q}^{\text{gr}}:\text{gr}^{pq}\widetilde{C}^{k}(\lambda,\mu)\to\text{gr}^{p\,q+1}\widetilde{C}^{k}(\lambda,\mu))}{\text{im}(\widetilde{Q}^{\text{gr}}:\text{gr}^{p\,q-1}\widetilde{C}^{k}(\lambda,\mu)\to\text{gr}^{pq}\widetilde{C}^{k}(\lambda,\mu))},$ (3.25)
where $\widetilde{Q}^{\text{gr}}$ is the differential on $\text{gr}(\widetilde{C}^{k}(\lambda,\mu))$ induced from $\widetilde{Q}.$
Some structural properties of the generalized affine $W$-algebra $W^{k}(\lambda,\mu)$ will be derived by using the cohomology of the graded algebra. The following two lemmas will be essential to derive the main result of this section. Recall the character $\chi$ as defined in (2.10) and the subspace $\mathfrak{p}$ defined in (3.19).
Lemma 3.8.
Extend the domain of the character $\chi$ to $\mathfrak{a}$ by letting $\chi(p)=0$ for any $p\in\mathfrak{p}.$ Then the map
$\varphi:\mathfrak{p}\to\mathfrak{a}^{\ast}\to\mathfrak{n}_{\lambda,\mu}^{\ast},\qquad a\mapsto a\cdot\chi\mapsto(a\cdot\chi)|_{\mathfrak{n}_{\lambda,\mu}}$ (3.26)
is surjective.
Proof.
If $E_{ij}^{(r)}\in\mathfrak{n}_{\lambda,\mu}$, then $\text{col}_{\mu}(i)<\text{col}_{\mu}(j)$ and $\lambda_{j}-\text{min}\{\lambda_{i},\lambda_{j}\}\leqslant r<\lambda_{j}$ and hence $E_{j-1,i}^{(\lambda_{j}-r-1)}$ is an element in $\mathfrak{p}.$ Observe that
$\varphi(E_{j-1,i}^{(\lambda_{j}-r-1)})(E_{i^{\prime},j^{\prime}}^{(r^{\prime})})=(E_{j-1,i}^{(\lambda_{j}-r-1)}\cdot\chi)(E_{i^{\prime},j^{\prime}}^{(r^{\prime})})=\chi([E_{i^{\prime}j^{\prime}}^{(r^{\prime})},E_{j-1,i}^{(\lambda_{j}-r-1)}])$
$=\delta_{j^{\prime},j-1}\delta_{i^{\prime},i-1}\delta_{r^{\prime},r+\lambda_{i}-\lambda_{j}}\delta_{\text{col}_{\mu}(i-1)+1,\text{col}_{\mu}(i)}-\delta_{i,i^{\prime}}\delta_{j,j^{\prime}}\delta_{r,r^{\prime}}\delta_{\text{col}_{\mu}(j-1)+1,\text{col}_{\mu}(j)}$
and so
$\varphi(E_{j-1,i}^{(\lambda_{j}-r-1)})=E_{j-1,i}^{(\lambda_{j}-r-1)}\cdot\chi=(E_{i-1,j-1}^{(\lambda_{i}-\lambda_{j}+r)})^{\ast}\delta_{\text{col}_{\mu}(i-1)+1,\text{col}_{\mu}(i)}-(E_{ij}^{(r)})^{\ast}.$ (3.27)
Here we noted that $\text{col}_{\mu}(j)\neq 1$ which implies $\text{col}_{\mu}(j-1)+1=\text{col}_{\mu}(j)$.
Now we show that any element $(E_{ij}^{(r)})^{*}$ of $\mathfrak{n}_{\lambda,\mu}^{*}$ is contained in $\text{Im}\,\varphi$ by the induction on $\text{col}_{\mu}(i)$. If $\text{col}_{\mu}(i)=1$ then
$\varphi(E_{j-1,i}^{(\lambda_{j}-r-1)})=-(E_{ij}^{(r)})^{\ast}$
by (3.27), i.e. $(E_{ij}^{(r)})^{\ast}\in\text{Im}(\varphi)$. Suppose $\text{col}_{\mu}(i)=c>1$. By the induction hypothesis, we have $(E_{i-1,j-1}^{(\lambda_{i}-\lambda_{j}+r)})^{\ast}\in\text{Im}\,\varphi$ and
$\varphi(E_{j-1,i}^{(\lambda_{j}-r-1)})=(E_{i-1,j-1}^{(\lambda_{i}-\lambda_{j}+r)})^{\ast}-(E_{ij}^{(r)})^{\ast}\in\text{Im}\,\varphi.$
Hence $(E_{ij}^{(r)})^{\ast}\in\text{Im}\,\varphi.$ Therefore $\varphi$ is surjective. ∎
The next lemma concerns the properties of the complex $(\widetilde{C}^{k}(\lambda,\mu),\widetilde{Q})$. To state it, recall that $r_{-}=\phi^{\mathfrak{n}_{\lambda,\mu}^{*}}\oplus J_{\mathfrak{p}}$ and set
$r_{-}^{pq}[\Delta]=\text{gr}^{pq}(\widetilde{C}^{k}(\lambda,\mu)[\Delta]\cap r_{-}).$
Lemma 3.9.
In the complex $(\widetilde{C}^{k}(\lambda,\mu),\widetilde{Q})$ we have the following.
  1. (1)
    For any $\Delta\in\mathbb{Z},$ the space $\widetilde{C}^{k}(\lambda,\mu)[\Delta]$ is finite dimensional.
  2. (2)
    We have the inclusion
    $\widetilde{Q}^{\text{gr}}(r_{-}^{pq}[\Delta])\subset r_{-}^{p\,q+1}[\Delta],$
    so that $\widetilde{Q}$ is almost a linear differential.
  3. (3)
    The differential $\widetilde{Q}$ on $\widetilde{C}^{k}(\lambda,\mu)$ is a good differential; that is,
    $H^{pq}(\text{gr}(\widetilde{C}^{k}(\lambda,\mu))[\Delta],\widetilde{Q}^{\text{gr}})=0$
    unless $p+q=0.$
Proof.
Part (1) is clear, while parts (2) and (3) follow directly from the fact that
$\widetilde{Q}^{\text{gr}}(J_{a})=-\phi^{a\cdot\chi}\qquad\text{and}\qquad\widetilde{Q}^{\text{gr}}(\phi^{\mathsf{m}})=0$
for $a\in\mathfrak{p}$ and $\mathsf{m}\in\mathfrak{n}^{\ast}_{\lambda,\mu}$, along with Lemma 3.8. ∎
Invoking [6, Lemma 4.18 and Theorem 4.19], we come to the following theorem.
Theorem 3.10.
Let $\{e_{i}\ |\ i\in S_{\chi}\}$ be a basis of $r_{-}^{p,-p}[\Delta]\cap\text{ker\,}\widetilde{Q}^{\text{gr}}$ and let
$E_{i}=e_{i}+\epsilon_{i}\in F^{p}(\widetilde{C}^{k}(\lambda,\mu)^{0})[\Delta]\cap\text{ker\,}\widetilde{Q}$
for some $\epsilon_{i}\in F^{p+\frac{1}{2}}\widetilde{C}^{k}(\lambda,\mu)^{0}[\Delta].$ Then
$H(R_{-},\widetilde{Q}):=\mathbb{C}[\partial]\otimes\text{Span}_{\mathbb{C}}\{E_{i}|i\in S_{\chi}\}$
is a nonlinear Lie conformal algebra. Moreover,
$H(\widetilde{C}^{k}(\lambda,\mu),\widetilde{Q})=H^{0}(\widetilde{C}^{k}(\lambda,\mu),\widetilde{Q})\simeq V(H(R_{-},\widetilde{Q}))$
where $V(H(R_{-},\widetilde{Q}))$ is the universal enveloping vertex algebra of $H(R_{-},\widetilde{Q}).$
As a consequence of Theorem 3.10, we get the following property of the generalized affine $W$-algebra:
$W^{k}(\lambda,\mu)\ \ {\text{\em is a vertex subalgebra of }}\ \ V(J_{\mathfrak{p}}),$ (3.28)
where $V(J_{\mathfrak{p}})$ is the universal enveloping vertex algebra of $\mathbb{C}[\partial]\otimes J_{\mathfrak{p}}$ endowed with the $\lambda$-bracket introduced in (3.20). Take an ordered basis $\mathcal{B}_{\mathfrak{p}}$ of $J_{\mathfrak{p}}$. Then $V(J_{\mathfrak{p}})$ has the PBW basis induced from the ordered set ~=$\widetilde{\mathcal{B}}=\{\,\partial^{t}\mathcal{B}_{\mathfrak{p}}\,|\,t\in\mathbb{Z}_{+}\,\}$ and we can define degrees of elements of $V(J_{\mathfrak{p}})$ by degrees with respect to the basis elements in $\widetilde{\mathcal{B}}$. In particular, the linear term of any element will mean the degree $1$ part in $V(J_{\mathfrak{p}})$ with respect to the given basis of $J_{\mathfrak{p}}$.
Corollary 3.11.
Let $a_{1},\dots,a_{r}$ be a basis of $\text{ker}\,\varphi$ for the map $\varphi$ in (3.26).
  1. (1)
    As a differential algebra, $W^{k}(\lambda,\mu)$ has a generating set consisting of $r=\text{dim}\,\mathfrak{a}(0)$ elements. Moreover, $r$ is the minimal number of elements in a generating set.
  2. (2)
    There exists a differential algebraically independent generating set $\{w_{i}\ |\ i=1,\dots,r\}$ of $W^{k}(\lambda,\mu)$ satisfying the following properties:
    1. (i)
      The element $w_{i}$ is homogeneous with respect to the conformal weight.
    2. (ii)
      The linear term of $w_{i}$ without total derivative is $J_{a_{i}}$, with respect to any basis of $J_{\mathfrak{p}}$.
    Moreover, any subset $\{w_{i}\ |\ i=1,\dots,r\}\subset W^{k}(\lambda,\mu)$ with properties (i) and (ii), is a generating set of $W^{k}(\lambda,\mu)$.
Proof.
By Theorem 3.10, the number of free generators of $W^{k}(\lambda,\mu)$ is dim(ker $\text{dim}(\text{ker }\varphi)$. Lemma 3.8 implies that
dim(ker $\text{dim}(\text{ker }\varphi)=\text{dim }\mathfrak{p}-\text{dim }\mathfrak{n}_{\lambda,\mu}^{\ast}=\text{dim }\mathfrak{p}-\text{dim }\mathfrak{n}_{\lambda,\mu}=\text{dim }\mathfrak{a}(0).$
Hence part (1) holds. Part (2) is a direct consequence of Theorem 3.10. ∎
Remark 3.12.
In the particular case $\lambda=(1^{N})$, our BRST complex defining the $W$-algebra $W^{k}(\lambda,\mu)$ coincides with the one used in [6, Sec. 5] to describe the $W$-algebra $W^{k}(\mathfrak{gl}_{N},f)$ associated with a nilpotent element $f\in\mathfrak{gl}_{N}$ of type $\mu$. Therefore, these two affine $W$-algebras are the same. ∎
4 Application of the Zhu functor
4.1 Zhu algebra of $W^{k}(\lambda,\mu)$
In this section, we obtain the Zhu algebra of generalized affine $W$-algebra which will be proved to be isomorphic to $U(\lambda,\mu).$ We start by recalling basic definitions related to the Zhu functor; see e.g. [6] for a more detailed discussion.
An operator $H$ on a vertex algebra $V$ is called a Hamiltonian if it is a diagonalizable operator satisfying the property:
$\Delta(a_{(n)}b)=\Delta(a)+\Delta(b)-n-1,$ (4.1)
where $a$ and $b$ are eigenvectors for $H$ and $H(c)=:\Delta(c)c$ for any eigenvector $c$. The $H$-twisted Zhu algebra $Zhu_{H}(V)$ of $V$ is the associative algebra given by
$Zhu_{H}(V)=V/(\partial+H)V.$
The associative algebra product and the commutator relation on $Zhu_{H}(V)$ are given by
$Zhu_{H}(a)\cdot Zhu_{H}(b)=Zhu_{H}(:ab:)+\int^{1}_{0}[Zhu_{H}(H(a)),Zhu_{H}(b)]_{x}dx,$ (4.2)
$[Zhu_{H}(a),Zhu_{H}(b)]:=[Zhu_{H}(a),Zhu_{H}(b)]_{\hbar=1}$
where
$[Zhu_{H}(a),Zhu_{H}(b)]_{\hbar}=\sum_{j\in\mathbb{Z}_{+}}{\Delta_{a}-1\choose j}\,\hbar^{j}\,Zhu_{H}(a_{(j)}b).$
Note that the conformal weight (3.21) on the complex $\widetilde{C}^{k}(\lambda,\mu)$ has the property (4.1). The conformal weight (3.21) can also be extended to $C^{k}(\lambda,\mu)$ by letting
$\Delta(a)=1-j_{a},\quad\Delta(\phi_{\mathsf{n}})=1-j_{\mathsf{n}}$ (4.3)
for $a\in\mathfrak{a}(j_{a}),$ and $\mathsf{n}\in\mathfrak{a}(j_{\mathsf{n}})$. Consider the Hamiltonian operator $H$ on $C^{k}(\lambda,\mu)$ defined by (4.3) and let
$C^{\text{fin}}(\lambda,\mu):=Zhu_{H}(C^{k}(\lambda,\mu)).$
We set
$\bar{a}:=Zhu_{H}(a)-k(h|a),\quad\bar{\phi}^{\mathsf{m}}:=Zhu_{H}(\phi^{\mathsf{m}}),\quad\bar{\phi}_{\mathsf{n}}:=Zhu_{H}(\phi_{\mathsf{n}}),\quad\bar{J_{a}}:=Zhu_{H}(J_{a}),$
where
$h=\sum_{\begin{subarray}{c}i=1,\dots,n\\ \text{col}_{\mu}(i)=r\end{subarray}}\ (n-r)\cdot E_{ii}^{(0)}.$
Note that $[h,a]=j_{a}a$ for $a\in\mathfrak{a}(j_{a})$. In other words, the grading $\bigoplus_{m\in\mathbb{Z}}\mathfrak{a}(m)$ can be induced from the Dynkin grading by $h$ on $\mathfrak{gl}_{N}$. Using (4.2), we obtain
$[\bar{a},\bar{b}]=[Zhu_{H}(a),Zhu_{H}(b)]=Zhu_{H}([a,b])+(\Delta_{a}-1)(a|b)k=Zhu_{H}([a,b])-j_{a}(a|b)k$
and hence
$[\bar{a},\bar{b}]=\overline{[a,b]},\quad[\bar{\phi}^{\mathsf{m}},\bar{\phi}_{\mathsf{n}}]=\mathsf{m}(\mathsf{n}).$
Let us consider the differential $d^{\text{fin}}:=Zhu_{H}(d)\in C^{\text{fin}}(\lambda,\mu)$ and $Q^{\text{fin}}=\text{ad }d^{\text{fin}}.$ Then we have
$Q^{\text{fin}}(\bar{a})$ $=Zhu_{H}(d_{(0)}a)$ (4.4)
$=Zhu_{H}(\phi^{I})Zhu_{H}([E_{I},a])-\sum_{I\in S_{\lambda,\mu}}k([h,E_{I}]|a)Zhu_{H}\phi^{I}=\sum_{I\in S_{\lambda,\mu}}\bar{\phi}^{I}\overline{[E_{I},a]},$
$Q^{\text{fin}}(\bar{\phi}_{\mathsf{n}})$ $=\bar{\mathsf{n}}+\chi(\mathsf{n})+\sum_{I\in S_{\lambda,\mu}}\bar{\phi}^{I}\bar{\phi}_{[E_{I},\mathsf{n}]},\ \ Q^{\text{fin}}(\bar{\phi}^{\mathsf{m}})=\frac{1}{2}\sum_{I\in S_{\lambda,\mu}}\bar{\phi}^{I}\bar{\phi}^{E_{I}\cdot\mathsf{m}},$
$Q^{\text{fin}}(\bar{J}_{a})$ $=\bar{\phi}^{I}\bar{J}_{\pi_{\leq}[E_{I},a]}-\bar{\phi}^{a\cdot\chi}+\bar{\phi}^{\pi_{\leq}[a,[h,E_{I}]]\cdot E_{I}^{\ast}}$
$-\sum_{I\in S_{\lambda,\mu}}\Big{(}k([h,E_{I}]|a)+\text{tr}_{\mathfrak{n}_{\lambda,\mu}}((\pi_{+}\text{ad }[h,E_{I}])\circ(\pi_{+}\text{ad }a))\Big{)}\bar{\phi}^{I}\,.$
Note that equality (4.4) holds because the conformal weight of the element $d$ is equal to $1$. Consider the decomposition
$C^{\text{fin}}(\lambda,\mu)=C^{\text{fin}}_{+}(\lambda,\mu)\otimes\widetilde{C}^{\text{fin}}(\lambda,\mu),$
where $C^{\text{fin}}_{+}(\lambda,\mu)$ and $\widetilde{C}^{\text{fin}}(\lambda,\mu)$ are subalgebras of $C^{\text{fin}}(\lambda,\mu)$ respectively generated by the subspaces $\bar{\phi}_{\mathfrak{n}_{\lambda,\mu}}\oplus\bar{J}_{\mathfrak{n}_{\lambda,\mu}}$ and $\bar{\phi}^{\mathfrak{n}_{\lambda,\mu}^{\ast}}\oplus\bar{J}_{\mathfrak{p}}$. Then the operator $H$ can be regarded as a Hamiltonian operator on both $C^{k}_{+}(\lambda,\mu)$ and $\widetilde{C}^{k}(\lambda,\mu)$ and
$Zhu_{H}(C^{k}_{+}(\lambda,\mu))=C^{\text{fin}}_{+}(\lambda,\mu),\quad Zhu_{H}(\widetilde{C}^{k}(\lambda,\mu))=\widetilde{C}^{\text{fin}}(\lambda,\mu).$
Similarly to the affine case, the differential $Q^{\text{fin}}$ satisfies
$Q^{\text{fin}}(C^{\text{fin}}_{+}(\lambda,\mu))\subset C^{\text{fin}}_{+}(\lambda,\mu),\qquad\bar{Q}(\widetilde{C}^{\text{fin}}(\lambda,\mu))\subset\widetilde{C}^{\text{fin}}(\lambda,\mu)$
and
$H(C^{\text{fin}}_{+}(\lambda,\mu),Q^{\text{fin}}|_{C^{\text{fin}}_{+}(\lambda,\mu)})=\mathbb{C}.$
Therefore,
$H(C^{\text{fin}}(\lambda,\mu),\bar{Q})\cong H(C^{\text{fin}}_{+}(\lambda,\mu),\bar{Q}|_{C^{\text{fin}}_{+}(\lambda,\mu)})\otimes H(\widetilde{C}^{\text{fin}}(\lambda,\mu),\widetilde{Q}^{\text{fin}})\cong H(\widetilde{C}^{\text{fin}}(\lambda,\mu),\widetilde{Q}^{\text{fin}}),$
where $\widetilde{Q}^{\text{fin}}:=Q^{\text{fin}}|_{\widetilde{C}^{\text{fin}}(\lambda,\mu)}$.
On the other hand, since the Hamiltonian operator $H$ on $C^{k}(\lambda,\mu)$ induces the Hamiltonian operator on the affine $W$-algebra $W^{k}(\lambda,\mu)$, we get the associative algebra $Zhu_{H}(W^{k}(\lambda,\mu)).$ By applying [6, Theorem 4.20], we come to the following proposition.
Proposition 4.1.
The following two associative algebras are isomorphic:
$Zhu_{H}(W^{k}(\lambda,\mu))\simeq H(C^{\text{\rm fin}}(\lambda,\mu),Q^{\text{\rm fin}}).$
Proof.
We have the grading (3.22) and filtration (3.23) on $\widetilde{C}^{k}(\lambda,\mu)$ satisfying
$F^{p_{1}}\widetilde{C}^{k}(\lambda,\mu)^{n_{1}}[\Delta_{1}]_{(n)}F^{p_{2}}\widetilde{C}^{k}(\lambda,\mu)^{n_{2}}[\Delta_{2}]\subset F^{p_{1}+p_{2}}\widetilde{C}^{k}(\lambda,\mu)^{n_{1}+n_{2}}[\Delta_{1}+\Delta_{2}-n-1]$
and
$\widetilde{Q}(F^{p}\widetilde{C}^{k}(\lambda,\mu)^{n}[\Delta])\subset F^{p}\widetilde{C}^{k}(\lambda,\mu)^{n+1}[\Delta].$
Furthermore, due to Lemma 3.9, the assumptions of [6, Theorem 4.20] are satisfied. Thus, there is a canonical associative algebra isomorphism
$Zhu_{H}(W^{k}(\lambda,\mu))=Zhu_{H}H(\widetilde{C}^{k}(\lambda,\mu),\widetilde{Q})\cong H(Zhu_{H}(\widetilde{C}^{k}(\lambda,\mu)),\widetilde{Q})=H(\widetilde{C}^{\text{fin}}(\lambda,\mu),\widetilde{Q}^{\text{fin}}),$
completing the proof. ∎
4.2 The algebras $U(\lambda,\mu)$ and $Zhu_{H}W^{k}(\lambda,\mu)$
We will show that the Zhu algebra of $W^{k}(\lambda,\mu)$ is isomorphic to the generalized finite $W$-algebra $U(\lambda,\mu)$ by analyzing the graded algebra of $U(\lambda,\mu)$. To this end, we need to understand $U(\lambda,\mu)$ via Lie algebra cohomology. As the first step, construct the classical finite version of BRST complex. Denote $S(V)$ the supersymmetric algebra of a vector superspace $V$ and set
$\mathcal{C}(\lambda,\mu):=S(\mathfrak{a})\otimes S(\phi_{\mathfrak{n}_{\lambda,\mu}}\oplus\phi^{\mathfrak{n}_{\lambda,\mu}^{\ast}}),$
where $\phi_{\mathfrak{n}_{\lambda,\mu}}$, $\phi^{\mathfrak{n}_{\lambda,\mu}^{\ast}}$ are the odd vector superspaces defined in (3.3). The Poisson bracket on $\mathcal{C}(\lambda,\mu)$ is induced from the Lie bracket on $\mathfrak{a}$ and the super-commutator $[\phi^{\mathsf{m}},\phi_{\mathsf{n}}]=\mathsf{m}(\mathsf{n}).$ Consider the element
$d^{cl}:=\sum_{I\in S_{\lambda,\mu}}\phi^{I}E_{I}+\phi^{\chi}+\frac{1}{2}\sum_{I,I^{\prime}\in S_{\lambda,\mu}}\phi^{I}\phi^{I^{\prime}}\phi_{[E_{I^{\prime}},E_{I}]}\in\mathcal{C}(\lambda,\mu)$
and set
$\mathcal{Q}:=\text{ad }d^{cl}=\{d^{cl},\cdot\}\in\text{End}\,\mathcal{C}(\lambda,\mu).$
By direct computation, one can show $\mathcal{Q}^{2}=0$ so that $\mathcal{Q}$ is a differential of the complex $\mathcal{C}(\lambda,\mu).$ For $a\in\mathfrak{a}$ set
$J^{cl}_{a}=a+\sum_{I\in S_{\lambda,\mu}}\phi^{I}\phi_{[E_{I},a]}\in\mathcal{C}(\lambda,\mu)$
and introduce the spaces
$J^{cl}_{\mathfrak{n}_{\lambda,\mu}}:=\{J^{cl}_{\mathsf{n}}|\mathsf{n}\in\mathfrak{n}_{\lambda,\mu}\},\qquad J^{cl}_{\mathfrak{p}}:=\{J^{cl}_{p}|p\in\mathfrak{p}\}.$
Then $\mathcal{C}_{+}(\lambda,\mu):=S(\phi_{\mathfrak{n}_{\lambda,\mu}}\oplus J^{cl}_{\mathsf{n}_{\lambda,\mu}})$ and $\widetilde{\mathcal{C}}(\lambda,\mu):=S(J^{cl}_{\mathfrak{p}}\oplus\phi^{\mathfrak{n}_{\lambda,\mu}^{\ast}})$ are Poisson subalgebras of $\mathcal{C}(\lambda,\mu)$. As in the affine case, we can show that
$\mathcal{Q}(\mathcal{C}_{+}(\lambda,\mu))\subset\mathcal{C}_{+}(\lambda,\mu),\quad\mathcal{Q}(\widetilde{\mathcal{C}}(\lambda,\mu))\subset\widetilde{\mathcal{C}}(\lambda,\mu).$
Moreover, since $\mathcal{C}(\lambda,\mu)\cong\mathcal{C}_{+}(\lambda,\mu)\otimes\widetilde{\mathcal{C}}(\lambda,\mu)$ and $H(\mathcal{C}_{+}(\lambda,\mu),\mathcal{Q}|_{\mathcal{C}_{+}(\lambda,\mu}))=\mathbb{C}$, we have
$H(\mathcal{C}(\lambda,\mu),\mathcal{Q})\cong H(\widetilde{\mathcal{C}}(\lambda,\mu),\widetilde{\mathcal{Q}}),$ (4.5)
where $\widetilde{\mathcal{Q}}:=\mathcal{Q}|_{\widetilde{\mathcal{C}}(\lambda,\mu)}.$ The differential $\widetilde{\mathcal{Q}}$ acts on the elements in $\widetilde{\mathcal{C}}(\lambda,\mu)$ as follows:
$\widetilde{\mathcal{Q}}(J^{cl}_{a})=\sum_{I\in S_{\lambda,\mu}}\phi^{I}J^{cl}_{\pi_{\leq}[E_{I},a]}-\phi^{a\cdot\chi},\qquad\widetilde{\mathcal{Q}}(\phi^{\mathsf{m}})=\frac{1}{2}\sum_{I\in S_{\lambda,\mu}}\phi^{I}\phi^{E_{I}\cdot\mathsf{m}}.$ (4.6)
Similar to the conformal weight on the complex $\widetilde{C}^{k}(\lambda,\mu)$, we define the $\Delta$-grading on $\widetilde{\mathcal{C}}(\lambda,\mu)$ and call it conformal weight, by setting
$\Delta(J^{cl}_{a})=1-j_{a},\qquad\Delta(\phi^{I})=\text{deg}(E_{I})$
for $a\in\mathfrak{a}(j_{a})$ and introduce the space
$\widetilde{\mathcal{C}}(\lambda,\mu)_{\Delta}=\text{Span}_{\mathbb{C}}\{A\in\widetilde{\mathcal{C}}(\lambda,\mu)|\Delta(A)\leqslant\Delta\}.$ (4.7)
We also consider the $\mathbb{Z}/2$-bigrading
$\text{gr}(J^{cl}_{a})=\left(j_{a}-\frac{1}{2},-j_{a}+\frac{1}{2}\right),\quad\text{gr}(\phi^{I})=\left(\text{deg}(E_{I})+\frac{1}{2},-\text{deg}(E_{I})+\frac{1}{2}\right)$
and the induced $\mathbb{Z}_{+}$-grading given by
$\widetilde{\mathcal{C}}(\lambda,\mu)=\bigoplus_{n\in\mathbb{Z}_{+}}\widetilde{\mathcal{C}}(\lambda,\mu)^{n},$
where
$\widetilde{\mathcal{C}}(\lambda,\mu)^{n}=\text{Span}_{\mathbb{C}}\big{\{}a_{1}a_{2}\dots a_{s}\ |\ a_{k}\in r_{-},\text{gr}(a_{k})=(p_{k},q_{k}),\sum_{k=1}^{s}(p_{k}+q_{k})=n\big{\}}$
for
$r_{-}=J^{cl}_{\mathfrak{p}}\oplus\phi^{\mathfrak{n}^{\ast}_{\lambda,\mu}}.$ (4.8)
Using the decreasing filtration
$F^{p}\widetilde{\mathcal{C}}(\lambda,\mu)=\text{Span}_{\mathbb{C}}\big{\{}a_{1}a_{2}\dots a_{s}\ |\ a_{k}\in r_{-},\text{gr}(a_{k})=(p_{k},q_{k}),\sum_{k=1}^{s}p_{k}\geqslant p\big{\}},$
define the associated graded algebra by
$\text{gr}(\widetilde{\mathcal{C}}(\lambda,\mu)):=\bigoplus_{p,q\in\mathbb{Z}/{2}}\text{gr}^{pq}\widetilde{\mathcal{C}}(\lambda,\mu)$
where
$\text{gr}^{pq}\widetilde{\mathcal{C}}(\lambda,\mu)=F^{p}\widetilde{\mathcal{C}}(\lambda,\mu)^{p+q})/F^{p+\frac{1}{2}}\widetilde{\mathcal{C}}(\lambda,\mu)^{p+q}$
and
$F^{p}\widetilde{\mathcal{C}}(\lambda,\mu)^{p+q}=F^{p}\widetilde{\mathcal{C}}(\lambda,\mu)\cap\widetilde{\mathcal{C}}(\lambda,\mu)^{p+q}.$
The corresponding graded cohomology is defined by
$H^{pq}(\text{gr}(\widetilde{\mathcal{C}}(\lambda,\mu)),\widetilde{\mathcal{Q}}^{\text{gr}})=\frac{\text{ker}(\widetilde{\mathcal{Q}}^{\text{gr}}:\text{gr}^{pq}\widetilde{\mathcal{C}}(\lambda,\mu)\to\text{gr}^{p\,q+1}\widetilde{\mathcal{C}}(\lambda,\mu))}{\text{im}(\widetilde{\mathcal{Q}}^{\text{gr}}:\text{gr}^{p\,q-1}\widetilde{\mathcal{C}}(\lambda,\mu)\to\text{gr}^{pq}\widetilde{\mathcal{C}}(\lambda,\mu))},$
where $\widetilde{\mathcal{Q}}^{\text{gr}}$ is the differential on $\text{gr}(\widetilde{\mathcal{C}}(\lambda,\mu))$ induced from $\widetilde{\mathcal{Q}}.$
Lemma 4.2.
The following properties hold.
  1. (1)
    The differential $\widetilde{\mathcal{Q}}$ on $\widetilde{\mathcal{C}}(\lambda,\mu)$ is good and $\text{\rm gr}^{pq}H(\widetilde{\mathcal{C}}(\lambda,\mu),\widetilde{\mathcal{Q}})\cong H^{pq}(\text{\rm gr}(\widetilde{\mathcal{C}}(\lambda,\mu)),\widetilde{\mathcal{Q}}^{\text{\rm gr}})$.
  2. (2)
    $H(\widetilde{\mathcal{C}}(\lambda,\mu),\widetilde{\mathcal{Q}})=H^{0}(\widetilde{\mathcal{C}}(\lambda,\mu),\widetilde{\mathcal{Q}})$.
Proof.
  1. (1)
    Similarly to Lemma 3.9 (3), we can show that $H^{pq}(\text{gr}(\widetilde{\mathcal{C}}(\lambda,\mu)),\widetilde{\mathcal{Q}}^{gr})=0$ unless $p+q=0$, which means $\mathcal{Q}$ is a good differential. Moreover, since $\mathcal{Q}$ preserves the conformal weight and the subspace of $\widetilde{C}(\lambda,\mu)$ with a given conformal weight is finite dimensional, we can apply [6, Lemma 4.2].
  2. (2)
    As in Lemma 3.9 (2), the differential $\widetilde{\mathcal{Q}}$ is almost linear. Moreover, recalling notation (4.8), observe that since $H^{pq}(r_{-},\widetilde{\mathcal{Q}}^{\text{gr}})$ is nonzero only when $p+q=0,$ the cohomology $H(\text{gr}(\widetilde{\mathcal{C}}(\lambda,\mu)),\widetilde{\mathcal{Q}}^{\text{gr}})\simeq S(H(r_{-},\widetilde{\mathcal{Q}}^{\text{gr}}))$ is also concentrated at degree 00 part. Now, by part (1) we get part (2).
Corollary 4.3.
The cohomology $H(\widetilde{\mathcal{C}}(\lambda,\mu),\widetilde{\mathcal{Q}})$ is a Poisson subalgebra of $S(\mathfrak{p}).$
Proof.
By the same argument as in the proof of Lemma 3.4, we can show $H(\mathcal{C}(\lambda,\mu),\mathcal{Q})$ has the Poisson algebra structure induced from that of $\mathcal{C}(\lambda,\mu)$. Since $H(\mathcal{C}(\lambda,\mu),\mathcal{Q})\simeq H^{0}(\widetilde{\mathcal{C}}(\lambda,\mu),\widetilde{\mathcal{Q}}),$ this is a Poisson subalgebra of $S(J_{\mathfrak{p}}^{cl}).$ Finally, since $J_{\mathfrak{p}}^{cl}\simeq\mathfrak{p}$ as Lie algebras, the proof is complete. ∎
Now we want to show that the Poisson algebra $H(\widetilde{\mathcal{C}}(\lambda,\mu),\widetilde{\mathcal{Q}})$ can also be realized in two different ways:
(1) Lie algebra cohomology of $\mathfrak{n}_{\lambda,\mu}$;   (2) graded algebra of $U(\lambda,\mu).$ (4.9)
Consider the Kazhdan filtration $K_{0}(\mathfrak{a})\subset K_{1}(\mathfrak{a})\subset K_{2}(\mathfrak{a})\subset\dots$ on $U(\mathfrak{a})$, where a homogeneous element $a\in\mathfrak{a}$ is in $K_{s}(\mathfrak{a})$ if and only if $a\in\mathfrak{a}(j_{a})$ for some $j_{a}\geqslant 1-s.$ Then for elements $a\in K_{s}(\mathfrak{a})$ and $b\in K_{t}(\mathfrak{a}),$ the commutator $ab-ba$ belongs to $K_{s+t-1}(\mathfrak{a}).$ Then the graded algebra $S(\mathfrak{a}):=\text{gr}^{K}(U(\mathfrak{a}))$ is the Poisson algebra endowed with the bracket
$\{\text{gr}^{K}_{s}(a),\text{gr}^{K}_{t}(b)\}=\text{gr}^{K}_{s+t-1}(ab-ba).$
In addition, the Kazhdan filtration induces the filtration on $U(\mathfrak{a})/\mathcal{I}_{\lambda,\mu}$ and
$\text{gr}^{K}(U(\mathfrak{a})/\mathcal{I}_{\lambda,\mu})=S(\mathfrak{a})/\mathcal{I}^{cl}_{\lambda,\mu},\qquad\mathcal{I}^{cl}_{\lambda,\mu}:=S(\mathfrak{a})\left<n+\chi(n)|n\in\mathfrak{n}_{\lambda,\mu}\right>.$
Then $U(\mathfrak{a})/\mathcal{I}_{\lambda,\mu}$ and $S(\mathfrak{a})/\mathcal{I}^{cl}_{\lambda,\mu}$ are $\mathfrak{n}_{\lambda,\mu}$-modules via adjoint actions. Here we note that the Kazhdan filtration is the analogue of the filtration (4.7) induced from the conformal weight. In the following proposition, we describe the two realizations of $H(\widetilde{\mathcal{C}}(\lambda,\mu),\widetilde{Q})$ in (4.9) more precisely via the $\mathfrak{n}_{\lambda,\mu}$-module $S(\mathfrak{a})/\mathcal{I}^{cl}_{\lambda,\mu}$.
Proposition 4.4.
The following two cohomologies are isomorphic:
$H(\widetilde{\mathcal{C}}(\lambda,\mu),\widetilde{\mathcal{Q}})\simeq H(\,\mathfrak{n}_{\lambda,\mu},S(\mathfrak{a})/\mathcal{I}^{cl}_{\lambda,\mu}\,)=H^{0}(\,\mathfrak{n}_{\lambda,\mu},S(\mathfrak{a})/\mathcal{I}^{cl}_{\lambda,\mu}\,).$
Hence $H(\widetilde{\mathcal{C}}(\lambda,\mu),\widetilde{\mathcal{Q}})\simeq(S(\mathfrak{a})/\mathcal{I}^{cl}_{\lambda,\mu})^{\text{ad}\,\mathfrak{n}_{\lambda,\mu}}.$
Proof.
We show that the Chevalley–Eilenberg complex of the $\mathfrak{n}_{\lambda,\mu}$-module $S(\mathfrak{a})/\mathcal{I}^{cl}_{\lambda,\mu}$ is isomorphic to the complex $(\widetilde{\mathcal{C}}(\lambda,\mu),\widetilde{\mathcal{Q}})$. Consequently, the cohomology of $(\widetilde{\mathcal{C}}(\lambda,\mu),\widetilde{\mathcal{Q}})$ is concentrated in degree 00, so the cohomology $H(\mathfrak{n}_{\lambda,\mu},S(\mathfrak{a})/\mathcal{I}^{cl}_{\lambda,\mu})$ is also concentrated in degree 00. By (4.6), the map between the complexes
$\wedge^{n}(\mathfrak{n}_{\lambda,\mu}^{\ast})\otimes S(\mathfrak{a})/\mathcal{I}^{cl}_{\lambda,\mu}$ $\to S(\phi^{\mathfrak{n}_{\lambda,\mu}^{\ast}}\oplus J^{cl}_{\mathfrak{p}}),$
$\mathsf{m}_{1}\wedge\mathsf{m}_{2}\wedge\ldots\wedge\mathsf{m}_{n}\otimes p$ $\mapsto\phi^{\mathsf{m}_{1}}\phi^{\mathsf{m}_{2}}\dots\phi^{\mathsf{m}_{n}}J^{cl}_{p}$
for $\mathsf{m}\in\mathfrak{n}^{*}_{\lambda,\mu}$ and $p\in\mathfrak{p}$ is an isomorphism. ∎
Now using Proposition 4.4, we can realize the generalized finite $W$-algebra $U(\lambda,\mu)$ via Lie algebra cohomology.
Proposition 4.5.
The Lie algebra cohomology of $U(\mathfrak{a})/\mathcal{I}_{\lambda,\mu}$ is concentrated on degree 00 part; that is,
$H(\,\mathfrak{n}_{\lambda,\mu},U(\mathfrak{a})/\mathcal{I}_{\lambda,\mu}\,)=H^{0}(\,\mathfrak{n}_{\lambda,\mu},U(\mathfrak{a})/\mathcal{I}_{\lambda,\mu}\,).$
Hence $U(\lambda,\mu)\simeq H^{0}(\,\mathfrak{n}_{\lambda,\mu},U(\mathfrak{a})/\mathcal{I}_{\lambda,\mu}\,)=H(\,\mathfrak{n}_{\lambda,\mu},U(\mathfrak{a})/\mathcal{I}_{\lambda,\mu}\,).$
Proof.
The main idea is the same as for the proof of [9, Proposition 5.2]. Consider the cochain complex
$0\longrightarrow U(\mathfrak{a})/\mathcal{I}_{\lambda,\mu}\longrightarrow\mathfrak{n}_{\lambda,\mu}^{\ast}\otimes U(\mathfrak{a})/\mathcal{I}_{\lambda,\mu}\longrightarrow\dots\longrightarrow\wedge^{n}\mathfrak{n}_{\lambda,\mu}^{\ast}\otimes U(\mathfrak{a})/\mathcal{I}_{\lambda,\mu}\longrightarrow\cdots$
and the filtration on $\wedge^{n}\mathfrak{n}_{\lambda,\mu}^{\ast}\otimes U(\mathfrak{a})/\mathcal{I}_{\lambda,\mu}$ given by
$F_{p}(\wedge^{n}\mathfrak{n}_{\lambda,\mu}^{\ast}\otimes U(\mathfrak{a})/\mathcal{I}_{\lambda,\mu})=\big{\{}(E_{I_{1}}^{\ast}\wedge E_{I_{2}}^{\ast}\wedge\ldots\wedge E_{I_{n}}^{\ast})\otimes v\ |\ \sum_{k=1}^{n}(\text{col}_{\mu}(j_{k})-\text{col}_{\mu}(i_{k}))+j\leqslant p\big{\}}$
for $v\in K_{j}U(\mathfrak{a})/\mathcal{I}_{\lambda,\mu}$ and $I_{k}=(i_{k},j_{k},r_{k})$. Considering the corresponding spectral sequence and its convergence, we get the proposition. The isomorphism $U(\lambda,\mu)\simeq H^{0}(\,\mathfrak{n}_{\lambda,\mu},U(\mathfrak{a})/\mathcal{I}_{\lambda,\mu}\,)$ follows directly from the definition of $U(\lambda,\mu).$
Finally, we can prove the following theorem which is the main result in this section.
Theorem 4.6.
We have an isomorphism of associative algebras:
$Zhu_{H}(W^{k}(\lambda,\mu))\cong U(\lambda,\mu).$
Proof.
By Proposition 4.1 and Proposition 4.5, we know that
$Zhu_{H}(W^{k}(\lambda,\mu))\cong H(C^{\text{fin}}(\lambda,\mu),Q^{\text{fin}}),\qquad U(\lambda,\mu)=H(\mathfrak{n}_{\lambda,\mu},U(\mathfrak{a})/\mathcal{I}_{\lambda,\mu}).$
Hence it suffices to show that the complexes $(C^{\text{fin}}(\lambda,\mu),Q^{\text{fin}})$ and $(\wedge^{\bullet}\mathfrak{n}_{\lambda,\mu}^{\ast}\otimes U(\mathfrak{a})/\mathcal{I}_{\lambda,\mu},d^{c})$ are quasi-isomorphic, where
$d^{c}(\Psi\otimes v)$ $=\frac{1}{2}\sum_{I\in S_{\lambda,\mu}}(E_{I})^{\ast}\wedge E_{I}\cdot\Psi\otimes v+\sum_{I\in S_{\lambda,\mu}}(E_{I})^{\ast}\wedge\Psi\otimes\text{ad}(E_{I})(v).$ (4.10)
Here we set $E_{I}=E_{ij}^{(r)}=0$ if $r\geqslant\lambda_{j}$ or $r<\lambda_{j}-\text{min}(\lambda_{i},\lambda_{j})$. Observe that
$U(\mathfrak{a})/\mathcal{I}_{\lambda,\mu}\simeq U(\mathfrak{a})\otimes_{\mathfrak{n}_{\lambda,\mu}}\mathbb{C}_{-\chi},$
where $\mathbb{C}_{-\chi}$ is the one-dimensional representation of $\mathfrak{n}_{\lambda,\mu}$ with $\mathsf{n}\cdot 1=-\chi(\mathsf{n})$. Hence we can follow the argument of [6, Appendix]. ∎
By Corollary 3.11 and Theorem 4.6, we get the following corollary.
Corollary 4.7.
Let $a_{1},\dots,a_{r}$ be a basis of $\text{ker}\,\varphi$ for the map $\varphi$ in (3.26).
  1. (1)
    As an associative algebra, $U(\lambda,\mu)$ has a generating set consisting of $r$ elements, where $r$ is the dimension of $\mathfrak{a}(0)$. Moreover, $r$ is the minimal number of elements in a generating set.
  2. (2)
    Suppose $a_{i}\in K_{\Delta_{i}}(\mathfrak{a})\setminus K_{\Delta_{i}-1}(\mathfrak{a})$ for $i=1,2,\dots,r.$ Let $v_{i}\in K_{\Delta_{i}}(\mathfrak{a})$ be an element in $U(\lambda,\mu)$ such that the linear term in $\text{gr}^{K}_{\Delta_{i}}(v_{i})$ is $a_{i}.$ Then the set $\{v_{i}\ |\ i=1,\dots,r\}$ generates $U(\lambda,\mu)$ and is algebraically independent.
5 Principal and minimal nilpotent $\mu$
Here we will discuss the $W$-algebras for two particular cases of $\mu$, while keeping $\lambda$ arbitrary. By extending the terminology from the column-partition $\lambda$, we will refer to the cases $\mu=(n)$ and $\mu=(1^{n-2}2)$ as the principal and minimal nilpotent cases, respectively; cf. Remark 3.12.
5.1 Principal nilpotent case
In this subsection, we describe generators of the generalized affine and finite $W$-algebras when $\mu=(n)$. In this particular case, the affine $W$-algebra $W^{k}(\lambda,(n))$ was introduced and described in [14], although the finite counterpart $U(\lambda,(n))$ was not discussed there. We will prove here that the center of $U(\mathfrak{a})$ is isomorphic to $U(\lambda,(n)).$
Keeping the definitions from Sec. 2, note that the dimension of $\mathfrak{a}(0)$ is the number of boxes in the pyramid $\lambda$, which is equal to $N$. Hence $\text{dim}(\text{ker}\,\varphi)=N$. We can decompose
$\text{ker}\,\varphi=\bigoplus_{m=1}^{n}(\text{ker}\,\varphi)[m]\subset\mathfrak{p}$
where $(\text{ker}\,\varphi)[m]$ is the subspace of $\text{ker}\,\varphi$ with conformal weight $m$. Then
$\text{dim}(\text{ker}\,\varphi)[m]=\lambda_{n-m+1}$ (5.1)
since the set
$E_{n,n-m+1}^{(r)}+E_{n-1,n-m}^{(r+\lambda_{n}-\lambda_{n-1})}+E_{n-2,n-m-1}^{(r+\lambda_{[n-1,n]}-\lambda_{[n-m,n-m+1]})}+\dots+E_{m,1}^{(r+\lambda_{[m+1,n]}-\lambda_{[2,n-m+1]})}$
with $10\leqslant r\leqslant\lambda_{n-m+1}-$ 1 is a basis of $(\text{ker}\,\varphi)[m].$ Combining Corollary 3.11 and (5.1), we can derive following properties.
Lemma 5.1.
Let $A:=\{w_{1},\dots,w_{N}\}$ be a subset of $W^{k}(\lambda,(n))$ and $m=1,2,\dots,n$. If $A$ has $\lambda_{n-m+1}$ elements whose linear parts without total derivatives span a subspace in $\mathfrak{p}$ of dimension $\lambda_{n-m+1}$ with conformal weight $m$, then $A$ freely generates $W^{k}(\lambda,(n))$ as a differential algebra.
Note that this agrees with [14], where generators of $W^{k}(\lambda,(n))$ were described.
Example 5.2.
Let $\lambda=(2,2)$. Recall from (3.28) that $W^{k}(\lambda,(n))$ can be regarded as a vertex subalgebra of $V(J_{\mathfrak{p}}).$ According to [14], conformal weight $1$ generators of $W^{k}(\lambda,(n))$ are
$w_{1}=J_{E_{11}^{(0)}}+J_{E_{22}^{(0)}},\quad w_{2}=J_{E_{11}^{(1)}}+J_{E_{22}^{(1)}}$
and conformal weight $2$ generators of $W^{k}(\lambda,(n))$ are
$w_{3}=J_{E_{21}^{(0)}}+:J_{E_{11}^{(0)}}J_{E_{22}^{(1)}}:+:J_{E_{11}^{(1)}}J_{E_{22}^{(0)}}:-(k+2)\partial J_{E_{11}^{(1)}},\quad w_{4}=J_{E_{21}^{(1)}}+:J_{E_{11}^{(1)}}J_{E_{22}^{(1)}}:\,.$
The $\lambda$-brackets between generators are
$[w_{3}{}_{\lambda}w_{3}]=(\partial+2\lambda)\left(-(k+4)w_{4}+\left(\frac{k}{4}+1\right):w_{2}w_{2}:\right)$
and all other $\lambda$-brackets are 00. Note that, in this example, we can see that $W^{k}(\lambda,(n))$ does not have a conformal vector; see Remark 3.7. If we look for a conformal vector of weight $2$ in the form
$w:=c_{1}\partial w_{1}+c_{2}\partial w_{2}+c_{11}:w_{1}w_{1}:+c_{12}:w_{1}w_{2}:+c_{22}:w_{2}w_{2}:+c_{3}w_{3}+c_{4}w_{4}$
and solve the equation
$[w{}_{\lambda}w]=(\partial+2\lambda)w+c\lambda^{3}$
then there will be no non-trivial solution $(c_{1},c_{2},c_{11},c_{12},c_{22},c_{3},c_{4})$.
Now we want to find a generating set of $U(\lambda,(n))$. Combining again Corollary 4.7 and (5.1), we get the following lemma.
Lemma 5.3.
Let $F:=\{v_{1},\dots,v_{N}\}$ be a subset of $U(\lambda,(n))$ and $m=1,2,\dots,n$. Suppose $F$ has $\lambda_{n-m+1}$ elements in $K_{m}(\mathfrak{a})\setminus K_{m-1}(\mathfrak{a})$ whose linear parts in the image by $\text{gr}^{K}_{m}$ span a subspace in $\mathfrak{p}$ of dimension $\lambda_{n-m+1}$, where $K_{\bullet}$ is the Kazhdan filtration. Then $F$ freely generates $U(\lambda,(n))$ as an associative algebra.
To find a generating set of $U(\lambda,(n))$, recall a result from [13]. Let
$\mathcal{M}:=\left[\begin{array}[]{cccc}x+(n-1)\lambda_{1}+\epsilon_{11}(u)&\epsilon_{12}(u)&\cdots&\epsilon_{1n}(u)\\ \epsilon_{21}(u)&x+(n-2)\lambda_{2}+\epsilon_{22}(u)&\cdots&\epsilon_{2n}(u)\\ \vdots&\vdots&\ddots&\vdots\\ \epsilon_{n1}(u)&\epsilon_{n2}(u)&\cdots&x+\epsilon_{nn}(u)\end{array}\right]$
be the matrix with entries in $\mathbb{C}[x]\otimes U(\mathfrak{a})[u]$, where
$\epsilon_{ij}(u)=\left\{\begin{array}[]{ll}E_{ij}^{(0)}+E_{ij}^{(1)}u+\dots+E_{ij}^{(\lambda_{j}-1)}u^{\lambda_{j}-1}&\text{ if }i\geqslant j,\\[3.99994pt] E_{ij}^{(\lambda_{j}-\lambda_{i})}u^{\lambda_{j}-\lambda_{i}}+\dots+E_{ij}^{(\lambda_{j}-1)}u^{\lambda_{j}-1}&\text{ if }i<j.\end{array}\right.$
Let $\Phi_{m}(u)=\sum_{r\in\mathbb{Z}_{+}}\Phi_{m}^{(r)}u^{r}$ for $\Phi_{m}^{(r)}\in U(\mathfrak{a})$, be the coefficients in the expansion of the column determinant of $\mathcal{M}$,
$\text{cdet}\,\mathcal{M}=x^{n}+\Phi_{1}(u)x^{n-1}+\dots+\Phi_{n}(u).$
Then the following is an algebraically independent generating set of the center of the algebra $U(\mathfrak{a})$ [13, Cor. 2.7]:
$\{\,\Phi_{m}^{(r)}\,|\,\lambda_{n-m+2}+\lambda_{n-m+3}+\dots+\lambda_{n}<r+m\leqslant\lambda_{n-m+1}+\lambda_{n-m+2}+\dots+\lambda_{n}\,\}.$ (5.2)
Theorem 5.4.
Let $m=1,2,\dots,n$ and $\Phi_{m}^{(r)}\in U(\mathfrak{a})$ be elements in (5.2). Let $\Psi_{m}^{(r)}$ be the image of $\Phi_{m}^{(r)}$ in $U(\mathfrak{a})/\mathcal{I}_{\lambda,\mu}.$ Then
$\{\,\Psi_{m}^{(r)}\,|\,\lambda_{n-m+2}+\lambda_{n-m+3}+\dots+\lambda_{n}<r+m\leqslant\lambda_{n-m+1}+\lambda_{n-m+2}+\dots+\lambda_{n}\,\}$ (5.3)
is an algebraically independent generating set of $U(\lambda,(n))$.
Proof.
It is obvious that $\Psi_{m}^{(r)}\in U(\lambda,(n))$ since $[\mathsf{n},\Phi_{m}^{(r)}]=0$ for any $\mathsf{n}\in\mathfrak{n}_{\lambda,(n)}.$ Also, it is not hard to check that $\Psi_{m}^{(r)}$ is nonzero element in the quotient
$\big{(}K_{m}(\mathfrak{a})\setminus K_{m-1}(\mathfrak{a})\big{)}\big{/}\mathcal{I}_{\lambda,\mu}\big{(}K_{m}(\mathfrak{a})\setminus K_{m-1}(\mathfrak{a})\big{)}.$
Hence $\text{gr}^{K}_{m}(\Psi_{m}^{(r)})$ is nontrivial. Set
$l_{m}(u):=\sum_{r\in\mathbb{Z}_{+}}l_{m}^{(r)}u^{r},$
where $l_{m}^{(r)}$ is the linear part of $\text{gr}^{K}_{m}(\Psi_{m}^{(r)}).$ By Lemma 5.3, it is enough to show that
$\{\,l_{m}^{(r)}\,|\,\lambda_{n-m+2}+\lambda_{n-m+3}+\dots+\lambda_{n}<r+m\leqslant\lambda_{n-m+1}+\lambda_{n-m+2}+\dots+\lambda_{n}\,\}$ (5.4)
is a linearly independent subset in $\mathfrak{p}.$ By direct computations, we get
$l_{m}(u)$ $=\epsilon_{n,n-m+1}(u)+\epsilon_{n-1,n-m+2}(u)u^{\lambda_{n}-\lambda_{n-m+1}}$ (5.5)
$+\epsilon_{n-2,n-m-1}(u)u^{\lambda_{n}+\lambda_{n-1}-\lambda_{n-m+1}-\lambda_{n-m}}+\dots+\epsilon_{m,1}(u)u^{\lambda_{[m+1,n]}-\lambda_{[2,n-m+1]}}$
for $\lambda_{[s,t]}:=\lambda_{s}+\lambda_{s+1}+\dots+\lambda_{t}$ and hence
$l_{m}^{(r)}=E_{n,n-m+1}^{(r)}+E_{n-1,n-m+2}^{(r-\lambda_{n}+\lambda_{n-m+1})}+\dots+E_{m1}^{(r-\lambda_{[m+1,n]}+\lambda_{[2,n-m+1]})}.$
This implies that the set (5.4) is linearly independent, completing the proof. ∎
Example 5.5.
Let $n=3.$ Let us verify that (5.3) is an algebraically independent generating set of $U(\lambda,(3)).$ In order to find the image of $\Psi_{m}^{(r)}$ in the graded algebra with respect to the Kazhdan filtration, consider the following matrix:
¯$\overline{\mathcal{M}}_{3}=\left[\begin{array}[]{ccc}x+2\lambda_{1}+\epsilon_{11}(u)&-u^{\lambda_{2}-1}&0\\ \epsilon_{21}(u)&x+\lambda_{2}+\epsilon_{22}(u)&-u^{\lambda_{3}-1}\\ \epsilon_{31}(u)&\epsilon_{32}(u)&x+\epsilon_{33}(u)\end{array}\right].$ (5.6)
We substituted $\epsilon_{12}(u)$, $\epsilon_{23}(u)$ and $\epsilon_{13}(u)$ by their respective images $-u^{\lambda_{2}-1}$, $-u^{\lambda_{3}-1}$ and 00 in $U(\mathfrak{a})/\mathcal{I}_{\lambda,\mu}.$ Since the graded algebra of $U(\mathfrak{a})/\mathcal{I}_{\lambda,\mu}$ is commutative, we can find $\text{gr}(\Psi_{m}^{(r)})$ by computing the column determinant of (5.6) which is given by
$\text{\rm cdet}\ \overline{\mathcal{M}}_{3}=x^{3}+x^{2}\Big{(}\epsilon_{11}+\epsilon_{22}+\epsilon_{33}+2\lambda_{1}+\lambda_{2}\Big{)}$
$+x\Big{(}(2\lambda_{1}+\epsilon_{11})(\lambda_{2}+\epsilon_{22})+(2\lambda_{1}+\epsilon_{11})\epsilon_{33}+(\lambda_{2}+\epsilon_{22})\epsilon_{33}+\epsilon_{32}u^{\lambda_{3}-1}+\epsilon_{21}u^{\lambda_{2}-1}\Big{)}$
$+\Big{(}(2\lambda_{1}+\epsilon_{11})(\lambda_{2}+\epsilon_{22})\epsilon_{33}+(2\lambda_{1}+\epsilon_{11})\epsilon_{32}u^{\lambda_{3}-1}+\epsilon_{21}\epsilon_{33}u^{\lambda_{2}-1}+\epsilon_{31}u^{\lambda_{2}+\lambda_{3}-2}\Big{)}.$
By setting $\Psi_{m}(u):=\sum_{r\in\mathbb{Z}_{+}}\Psi_{m}^{(r)}u^{r}$ and $\text{gr}\,\Psi_{m}(u):=\sum_{r\in\mathbb{Z}_{+}}\text{gr}(\Psi_{m}^{(r)})u^{r}$ we find
$\text{gr}\,\Psi_{1}(u)=\epsilon_{11}+\epsilon_{22}+\epsilon_{33}+2\lambda_{1}+\lambda_{2},$
$\text{gr}\,\Psi_{2}(u)=(2\lambda_{1}+\epsilon_{11})(\lambda_{2}+\epsilon_{22})+(2\lambda_{1}+\epsilon_{11})\epsilon_{33}+(\lambda_{2}+\epsilon_{22})\epsilon_{33}+\epsilon_{32}u^{\lambda_{3}-1}+\epsilon_{21}u^{\lambda_{2}-1},$
$\text{gr}\,\Psi_{3}(u)=(2\lambda_{1}+\epsilon_{11})(\lambda_{2}+\epsilon_{22})\epsilon_{33}+(2\lambda_{1}+\epsilon_{11})\epsilon_{32}u^{\lambda_{3}-1}+\epsilon_{21}\epsilon_{33}u^{\lambda_{2}-1}+\epsilon_{31}u^{\lambda_{2}+\lambda_{3}-2}.$
Hence $l_{m}(u)$ in (5.5) for $m=1,2,3$ are
$l_{1}=\epsilon_{11}+\epsilon_{22}+\epsilon_{33},\quad l_{2}=\epsilon_{32}+\epsilon_{21}u^{\lambda_{3}-\lambda_{2}},\quad l_{3}=\epsilon_{31}\ .$
Now, by direct computations, one can check that $l_{1}^{(r_{1})}$, $l_{2}^{(r_{2})}$, $l_{3}^{(r_{3})}$ for $r_{1}\in\{0,1,\dots,\lambda_{3}-1\}$, $r_{2}\in\{\lambda_{3}-1,\dots,\lambda_{2}+\lambda_{3}-2\}$, $r_{3}\in\{\lambda_{2}+\lambda_{3}-2,\dots,\lambda_{1}+\lambda_{2}+\lambda_{3}-3\}$ are linearly independent.
As a direct consequence of Theorem 5.4, we get the following corollary.
Corollary 5.6.
The generalized finite $W$-algebra $U(\lambda,(n))$ is isomorphic to the center of $U(\mathfrak{a})$ as an associative algebra. In particular, the algebra $U(\lambda,(n))$ is commutative.
5.2 Minimal nilpotent case
In this subsection, we describe generators of generalized finite and affine $W$-algebras when $\lambda$ is arbitrary and $\mu=(1,1,\dots,1,2)$ (with $n-2$ parts equal to $1$). From the $\mathbb{Z}$-grading (2.6),
$\mathfrak{n}_{\lambda,\mu}=\text{Span}_{\mathbb{C}}\{E_{\beta n}^{(r)}\ |\ 1\leqslant\beta\leqslant n-1,\lambda_{n}-\lambda_{\beta}\leqslant r<\lambda_{n}\}.$
To apply Corollaries 3.11 and 4.7, we first describe the kernel of $\varphi$. Since $\chi=(E_{n-1,n}^{(\lambda_{n}-1)})^{\ast}$, we have
$\text{ker}\,\varphi=\text{Span}_{\mathbb{C}}\left(\mathcal{B}_{1}\cup\mathcal{B}_{2}\right)$
where
$\mathcal{B}_{1}$ :=$:=\{E_{ij}^{(r)}\ |\ 1\leqslant i\leqslant n-2,1\leqslant j\leqslant n-1,\lambda_{j}-\text{min}(\lambda_{i},\lambda_{j})\leqslant r<\lambda_{j}\}$
[…]
Therefore the number of free generators of the generalized finite and affine $W$-algebras is
$\text{dim}\,\mathfrak{a}(0)=\text{dim}\,(\text{ker}\,\varphi)=(2n-3)\lambda_{1}+(2n-5)\lambda_{2}+\dots+\lambda_{n-1}+\lambda_{n}\ .$
Moreover, the numbers of generators of conformal weights $1$ and $2$ are
$|S_{1}|=(2n-4)\lambda_{1}+(2n-6)\lambda_{2}+\dots+2\lambda_{n-2}+\lambda_{n}\qquad\text{and}\qquad|S_{2}|=\lambda_{1}+\lambda_{2}+\dots+\lambda_{n-1},$
respectively. In the following theorems, we find generating sets of $U(\lambda,\mu)$ and $W^{k}(\lambda,\mu).$
Theorem 5.7.
The generalized finite $W$-algebra $U(\lambda,\mu)$ has the following properties:
  1. (1)
    Any element $a\in\mathcal{B}_{1}$ is in $U(\lambda,\mu)$ and has conformal weight 1.
  2. (2)
    For $E_{n\alpha}^{(r)}\in\mathcal{B}_{2}$, the element
    $E_{n\alpha}^{(r)}+\sum_{a+b=\lambda_{n}-1+r}E_{n-1,\alpha}^{(a)}E_{nn}^{(b)}-\sum_{\gamma=1}^{n-2}\ \sum_{a+b=\lambda_{n}-1+r}E_{n-1,\gamma}^{(a)}E_{\gamma\alpha}^{(b)}-\delta_{r,0}\delta_{\lambda_{\alpha},\lambda_{n}}\lambda_{n}E_{n-1,\alpha}^{(\lambda_{n}-1)}$ (5.7)
    is in $U(\lambda,\mu)$ and has conformal weight $2$.
  3. (3)
    The set $\mathcal{B}_{1}$ and the elements in (5.7) comprise an algebraically independent generating set of $U(\lambda,\mu).$
The proof of Theorem 5.7 is a simple analogue of the proof of the next theorem. Recall the property of $W^{k}(\lambda,\mu)$, established in (3.28).
Theorem 5.8.
For the generators of $W^{k}(\lambda,\mu),$ we have the following properties.
  1. (1)
    For $a\in\mathcal{B}_{1}$, the element $J_{a}$ is in $W^{k}(\lambda,\mu)$ and has conformal weight 1.
  2. (2)
    For $E_{n\alpha}^{(r)}\in\mathcal{B}_{2}$, the element
    $G_{n\alpha}^{(r)}:=J_{E_{n\alpha}^{(r)}}$ $+\sum_{a+b=\lambda_{n}-1+r}:J_{E_{n-1,\alpha}^{(a)}}J_{E_{nn}^{(b)}}:-\sum_{\gamma=1}^{n-2}\ \sum_{a+b=\lambda_{n}-1+r}:J_{E_{n-1,\gamma}^{(a)}}J_{E_{\gamma\alpha}^{(b)}}:$ (5.8)
    $-\delta_{r,0}\delta_{\lambda_{\alpha},\lambda_{n}}\left(\lambda_{n}+k\right)\partial J_{E_{n-1,\alpha}^{(\lambda_{n}-1)}}.$
    is in $W^{k}(\lambda,\mu)$ and has conformal weight $2$.
  3. (3)
    The set $\{J_{a}\ |\ a\in\mathcal{B}_{1}\}$ and elements in (5.8) comprise a differential algebraically independent generating set of $W^{k}(\lambda,\mu)$.
Proof.
  1. (1)
    By Theorem 3.10, it is sufficient to show that the above elements are inside $\widetilde{Q}(J_{a})=0$ for $a\in\mathcal{B}_{1}$. Let $21\leqslant i\leqslant n-$ 2 and $11\leqslant j\leqslant n-$ 1. Then
    $\widetilde{Q}(J_{E_{ij}^{(r)}})=-\delta_{jn}\phi^{E_{n-1,i}^{(\lambda_{n}-1-r)}{}^{\ast}}+\delta_{i,n-1}\phi^{E_{jn}^{(\lambda_{n}-1-r)}{}^{\ast}}=0$
    and
    $\widetilde{Q}(J_{E_{n-1,n-1}^{(r)}+E_{nn}^{(r)}})=-\phi^{E_{n-1,n}^{(\lambda_{n}-1-r)}{}^{\ast}}+\phi^{E_{n-1,n}^{(\lambda_{n}-1-r)}{}^{\ast}}=0\,.$
    Hence we get the assertion.
  2. (2)
    Now suppose $E_{n\alpha}^{(r)}\in\mathcal{B}_{2}$. Then
    $\widetilde{Q}(J_{E_{n\alpha}^{(r)}})$ $=\sum_{\gamma=1}^{n-1}\sum_{a,b}:\phi^{E_{\gamma n}^{(\lambda_{n}-1-a)}}{}^{\ast}J_{E_{\gamma\alpha}^{(b)}}:-\sum_{a,b}:\phi^{E_{\alpha n}^{(\lambda_{n}-1-a)}}{}^{\ast}J_{E_{nn}^{(b)}}:$
    $\qquad+\delta_{\lambda_{\alpha},\lambda_{n}}\delta_{r,0}\frac{k}{N}(\lambda_{1}+\dots+\lambda_{\alpha-1}+(n-\alpha+1)\lambda_{\alpha})\partial\phi^{E_{\alpha n}^{(0)}{}^{\ast}}$
    which equals
    $\phantom{=}\sum_{\gamma=1}^{n-1}\sum_{a,b}:\phi^{E_{\gamma n}^{(\lambda_{n}-1-a)}}{}^{\ast}J_{E_{\gamma\alpha}^{(b)}}:-\sum_{a,b}:\phi^{E_{\alpha n}^{(\lambda_{n}-1-a)}}{}^{\ast}J_{E_{nn}^{(b)}}:$
    $\qquad+\delta_{\lambda_{\alpha},\lambda_{n}}\delta_{r,0}\frac{k}{N}(\lambda_{1}+\dots+\lambda_{\alpha-1}+\lambda_{\alpha}+\lambda_{\alpha+1}+\dots+\lambda_{n})\partial\phi^{E_{\alpha n}^{(0)}}{}^{\ast}$
    $=\sum_{\gamma=1}^{n-1}\sum_{a,b}:\phi^{E_{\gamma n}^{(\lambda_{n}-1-a)}}{}^{\ast}J_{E_{\gamma\alpha}^{(b)}}:-\sum_{a,b}:\phi^{E_{\alpha n}^{(\lambda_{n}-1-a)}}{}^{\ast}J_{E_{nn}^{(b)}}:+\delta_{\lambda_{\alpha},\lambda_{n}}\delta_{r,0}k\partial\phi^{E_{\alpha n}^{(0)}}{}^{\ast},$
    where all the summation for $a,b$ runs over the all the integer values satisfying $a+b=\lambda_{n}-1+r.$ One the other hand, when $a+b=\lambda_{n}-1+r$ and $\gamma=1,\dots,n-2$, we have
    $\widetilde{Q}(:J_{E_{n-1,\alpha}^{(a)}}J_{E_{nn}^{(b)}}:)=:\phi^{E_{\alpha n}^{(\lambda_{n}-1-a)}{}^{\ast}}J_{E_{nn}^{(b)}}:-:\phi^{E_{n-1,n}^{(\lambda_{n}-1-a)}}{}^{\ast}J_{E_{n-1,\alpha}^{(b)}}:+\delta_{\lambda_{\alpha},\lambda_{n}}\delta_{r,0}\lambda_{n}\partial\phi^{E_{\alpha n}^{(0)}{}^{\ast}},$
    $\widetilde{Q}(:J_{E_{n-1,\gamma}^{(a)}}J_{E_{\gamma\alpha}^{(b)}}:)=:\phi^{E_{\gamma n}^{(\lambda_{n}-1-a)}}{}^{\ast}J_{E_{\gamma\alpha}^{(b)}}:,$
    $\widetilde{Q}(\partial J_{E_{n-1,\alpha}^{(\lambda_{n}-1)}})=\delta_{r,0}\delta_{\lambda_{n},\lambda_{\alpha}}\partial\phi^{E_{\alpha n}^{(0)}{}^{\ast}}.$
    These relations imply that $\widetilde{Q}(G_{n\alpha}^{(r)})=0$ and so this element is in $W^{k}(\lambda,\mu).$
  3. (3)
    This follows directly from Corollary 3.11.
Example 5.9.
Let $\lambda=(1,1,2,2)$ and $\mu=(1,1,2)$. Then
$\mathcal{B}_{1}$ =$=\{E_{11}^{(0)},E_{12}^{(0)},E_{13}^{(1)},E_{21}^{(0)},E_{22}^{(0)},E_{23}^{(1)}\}\cup\{E_{33}^{(0)}+E_{44}^{(0)},E_{33}^{(1)}+E_{44}^{(1)}\}\ ,$
$\mathcal{B}_{2}$ =$=\{E_{41}^{(0)},E_{42}^{(0)},E_{43}^{(1)},E_{43}^{(0)}\}\ .$
For $a\in\mathcal{B}_{1}$, $a$ and $J_{a}$ are generators of $U(\lambda,\mu)$ and $W^{k}(\lambda,\mu)$, respectively, of conformal weight $1$. The generators of $U(\lambda,\mu)$ of conformal weight $2$ are
$E_{41}^{(0)}+E_{31}^{(0)}E_{44}^{(1)}\ ,\quad E_{42}^{(0)}+E_{32}^{(0)}E_{44}^{(1)}\ ,\quad E_{43}^{(1)}+E_{33}^{(1)}E_{44}^{(1)}\ ,$
$E_{43}^{(0)}+E_{33}^{(0)}E_{44}^{(1)}+E_{33}^{(1)}E_{44}^{(0)}-E_{31}^{(0)}E_{13}^{(1)}-E_{32}^{(0)}E_{23}^{(1)}-2E_{33}^{(1)}$
and the generators of $W^{k}(\lambda,\mu)$ of conformal weight $2$ are
$J_{E_{41}^{(0)}}+:J_{E_{31}^{(0)}}J_{E_{44}^{(1)}}:\ ,\quad J_{E_{42}^{(0)}}+:J_{E_{32}^{(0)}}J_{E_{44}^{(1)}}:\ ,\quad J_{E_{43}^{(1)}}+:J_{E_{33}^{(1)}}J_{E_{44}^{(1)}}:\ ,$
$J_{E_{43}^{(0)}}+:J_{E_{33}^{(0)}}J_{E_{44}^{(1)}}:+:J_{E_{33}^{(1)}}J_{E_{44}^{(0)}}:-:J_{E_{31}^{(0)}}J_{E_{13}^{(1)}}:-:J_{E_{32}^{(0)}}J_{E_{23}^{(1)}}:-(k+2)\partial J_{E_{33}^{(1)}}\ .$
Example 5.10.
If $\lambda=(1,1,2,3)$ and $\mu=(1,1,2)$, then the last terms of (5.7) and (5.8) are equal to 00. In this case, the generators of $W^{k}(\lambda,\mu)$ of conformal weight $2$ are
$J_{E_{41}^{(0)}}$ $+:J_{E_{31}^{(0)}}J_{E_{44}^{(2)}}:\ ,\quad J_{E_{42}^{(0)}}+:J_{E_{32}^{(0)}}J_{E_{44}^{(2)}}:$
$J_{E_{43}^{(1)}}$ $+:J_{E_{33}^{(1)}}J_{E_{44}^{(2)}}:\ ,\quad J_{E_{43}^{(0)}}+:J_{E_{33}^{(0)}}J_{E_{44}^{(2)}}:\ .$
Declarations
Competing interests
The authors have no competing interests to declare that are relevant to the content of this article.
Acknowledgements
This project was initiated during the second named author visit to the Seoul National University. He is grateful to the Department of Mathematical Sciences for the warm hospitality. His work was also supported by the Australian Research Council, grant DP240101572. The work of U.R. Suh was supported by NRF Grant #2022R1C1C1008698 and Creative-Pioneering Researchers Program by Seoul National University.
Availability of data and materials
No data was used for the research described in the article.